### v2

# Advanced safety filter based on SOS Control Barrier and Lyapunov Functions

Abstract
This paper presents a novel safety filter framework that ensures both safety and the preservation of the legacy control action within a nominal region. This modular design allows the safety filter to be integrated into the control hierarchy without compromising the performance of the existing legacy controller within the nominal region. This is accomplished by formulating multiple Control Barrier Functions (CBFs) and Control Lyapunov-like Functions (CLFs) conditions, alongside a forward invariance condition for the legacy controller, as sum-of-squares constraints utilizing Putinar’s Positivstellensatz. Additionally, the state-dependent inequality constraints of the resulting Quadratic Program – encoding the CBF and CLF conditions – are designed to remain inactive within the nominal region, ensuring perfect tracking of the legacy control action. Our safety filter design is also the first to include quadratic input constraints, and does not need an explicit specification of the attractor, as it is implicitly defined by the legacy controller. To avoid the chattering effect and guarantee the uniqueness and Lipschitz continuity of solutions, the state-dependent inequality constraints of the Quadratic Program are selected to be sufficiently regular. Finally, we demonstrate the method in a detailed case study involving the control of a three-phase ac/dc power converter.
I Introduction
In various engineering applications, it is of much interest to endow a well-performant controller with safety guarantees by minimally changing its control action. This concept is commonly referred to as a safety filter, and a comprehensive overview of the subject can be found in [1]. An illustrative example is a finely-tuned control design based on linearization, which must be extended to adhere to hard state constraints in the nonlinear closed-loop system. We refer to such a well-performant controller as the legacy controller. A popular type of safety filters relies on solving a Quadratic Program (QP) at every state using a Control Barrier Function (CBF). By continuously solving the QP, the control action of the legacy controller is smoothly adjusted as the system approaches the boundary of the safe set $\mathcal{X}_{s}$. The main challenges in deploying such a safety filter lies in finding a viable CBF candidate and adhering to input constraints.
Even though safety filters based on a QP are well-established, they are frequently not adopted in current industrial setups due to their risk of altering the legacy control behavior during nominal operations. In practice, the legacy control parameters are often meticulously tuned and thoroughly evaluated, typically involving extensive hardware-in-the-loop simulations, which require significant time and resources. As a result, the safety filter is expected to intervene only when the system approaches state constraints, meaning when it is leaving the nominal operating regime.
Figure 1: The safe set $\mathcal{X}_{s}$, defined as the zero sublevel set of the CBF $B(x)$, specifies the set of states for which safety can be guaranteed. The safe set must be contained within the allowable set $\mathcal{X}_{a}$, which encodes the system’s state constraints. Finally, the nominal region $\mathcal{X}_{n}$, defined as the zero sublevel set of the CLF $V(x)$, must be contained within the safe set $\mathcal{X}_{s}$.
This paper presents a safety filter concept that guarantees the preservation of the legacy control action within a nominal region $\mathcal{X}_{n}$, which contains the desired closed-loop equilibrium point (or another attractor), as depicted in Figure 1. Unlike in a reach-avoid problem [2], the nominal region $\mathcal{X}_{n}$ does not represent a predetermined set of target states. Instead, it is selected during the design process (often at an earlier stage) to encompass the nominal operating regime of the legacy controller while maintaining a sufficiently large transitional region for a smooth safety filter operation. The specifications are inherently more stringent than those of a standard safety filter, but may therefore be deployed in an industrial setup without interference in the original line-up.
Our safety filter is realized based on solving a Quadratic Constraint Quadratic Program (QCQP) at every state using multiple CBFs and Control Lyapunov-like Functions (CLFs). A QCQP instead of a simple QP is employed to encode quadratic input constraints, which are common, e.g., in power electronic applications. While CBF conditions (encoded as linear constraints of the QCQP) are used to ensure safety w.r.t. $\mathcal{X}_{s}$, we leverage the CLF condition (also encoded as linear constraints of the QCQP) to specify a dissipation rate ensuring a finite-time convergence from the boundary of the safe set to the nominal region $\mathcal{X}_{n}$. The finite-time convergence property is not intended to enhance control performance but to ensure a rapid return to the nominal operating regime following an unforeseen event. We ensure that the nominal set contains the desired attractor by selecting it forward invariant with respect to the legacy controller. For identifying compatible CBF and CLF candidates that adhere to the linear and quadratic input constraints, as well as a solution for the QCQP at every state, we propose a Sum-of-Squares (SOS) optimization approach. Preserving the legacy control action is achieved by ensuring that the inequality constraints of the QCQP remain inactive within $\mathcal{X}_{n}$. This necessitates taking into account the legacy controller when searching for appropriate CBF and CLF candidates. To solve the resulting SOS program, we deploy an alternating algorithm that simultaneously searches for a feasible controller in the class of rational functions of the state.
The concept of our safety filter resembles the integration of a CLF with a CBF to attain both safety and stabilization. Such a unification, although just with a single function, has been initially proposed in [3]. The integration of both CLF and CBF into an optimization-based controller has since become a well-established concept, see the survey in [4]. Feasibility of such an optimization-based controller can be accomplished by introducing a slack variable that deactivates CLFs and CBFs based on their respective priorities, as shown in [5]. Alternatively, it is possible to compute compatible CLFs and CBFs using SOS tools, ensuring that for every state, there exists an input $u$ satisfying all CBF and CLF conditions. In order to encode compatibility, there is one approach that introduces an SOS constraint for each CLF/CBF pair in the absence of inputs constraints [6]. The second approach establishes compatibility by introducing a polynomial or rational state-feedback controller, as outlined in [7, 8]. Such a controller can also be used to enforce linear input constraints, as demonstrated in [9]. Another noteworthy approach for imposing input constraints using a contrapositive statement, albeit without ensuring compatibility between CBFs and CLFs, is introduced in [10]. These approaches introduce auxiliary polynomials that render the resulting SOS problem bilinear in its decision variables. In [11], an alternating algorithm is proposed to solve this bilinear SOS problem. One challenge in seeking CLF and CBF candidates through an alternating algorithm is to identify an initial feasible point. In their work in [12], the authors constrained their search for an initial CBF to scalar functions of a specific form. This restriction allows to reformulate the problem with linearity in its decision variables.
After our previous work [8] on synthesizing a rational controller to find compatible CLF and multiple CBFs using SOS optimization, and through collaboration with industrial partners, we identified a recurring issue: any modifications to the existing legacy control architecture should not compromise the nominal control performance. A prime example for this is the integration of a safety filter. In this paper, we propose a novel way of implementing a safety filter without disturbing the legacy controller within the nominal region. The key advancements are summarized as follows:
  • First, we propose a novel safety filter implementation that, by design, ensures finite-time convergence to the nominal region $\mathcal{X}_{n}$ and perfect tracking of the legacy control action within that region.
  • Second, multiple CBFs are used to the increase flexibility of the resulting safe set. Instead of imposing each CBF condition separately, we relax these conditions when the corresponding CBF is not active, similar to the idea in [13]. Moreover, the CLF condition can also be relaxed by leveraging the forward invariance property of the CBFs, similar to [14] for autonomous systems.
  • Third, the desired attractor (e.g., an equilibrium point or limit cycle), which is stabilized by the legacy controller in the nominal region, does not need to be specified explicitly, as it is implicitly specified through the legacy controller as a forward invariance condition.
  • Fourth, we introduce a novel approach for encoding linear and quadratic input constraints on the rational controller. This ensures feasibility of the resulting QCQP-based controller, even in case of quadratic input constraints – an issue that, to the authors’ knowledge, is unsolved.
To avoid chattering and guarantee the uniqueness and Lipschitz continuity of solutions, the state-dependent inequality constraints of the QCQP are selected to be sufficiently regular. Finally, we tested our framework on a detailed three-phase ac/dc converter system connected to an infinite bus, where we successfully synthesized a safety filter that ensures satisfaction of the state constraints given by the dc-link voltage and line current, while adhering to quadratic input constraints. We further tracked an a priori unspecified quadrature current reference of the ac/dc converter by integrating the reference as a state into a slightly modified system model.
The remainder of the paper is structured as follows: In Section II, the preliminaries and notation are discussed. Section III introduces the concept of the advanced safety filter, its properties, and the resulting problem statement. Numerical simulations, presented in Section VI, are provided to demonstrate the efficacy of our safety filter. Finally, Section VII concludes the paper.
II Preliminaries & Notation
II-A Notation
The shorthand $\mathcal{I}_{t}:=\{1,2,...,t\}$ designates a range of natural numbers from 1 to $t$. $R[x]$ refers to the set of scalar polynomials in variables $x\in\mathbb{R}^{n}$, and $\Sigma[x]$ refers to the set of scalar SOS polynomials in variables $x\in\mathbb{R}^{n}$. If $p(x)\in\Sigma[x]$, it can be expanded to
$p(x)=Z(x)^{T}QZ(x),$ (1)
where $Z(x)$ is a vector of monomials in $x$, and $Q$ is a square positive semidefinite matrix.
Throughout the paper, we consider a polynomial control system, affine in the control action $u\in\mathbb{R}^{m}$, and described as
$\dot{x}=F_{c}(x,u)=f(x)+G(x)u,$ (2)
where $x\in\mathbb{R}^{n}$ is the state, $f(x)\in\left(R[x]\right)^{n}$ is a polynomial vector field, and $G(x)\in\left(R[x]\right)^{n\times m}$ is a polynomial matrix. System 2 with a polynomial state feedback control policy $u_{cl}(x)\in\left(R[x]\right)^{m}$ results in a closed-loop system of the form
$\dot{x}=F_{cl}(x)=f(x)+G(x)u_{cl}(x).$ (3)
The interior of a set $\mathcal{X}\subseteq\mathbb{R}^{n}$ is denoted by $\text{int}(\mathcal{X})$. A subset $\mathcal{X}\subseteq\mathbb{R}^{n}$ is called forward invariant (cf. [15, Theorem 4.4]) with respect to system 3 if for every $x(0)\in\mathcal{X}$, $x(t)\in\mathcal{X}$ for all $t\geq 0$. A system 3 is called safe (cf. [16]) w.r.t. an allowable set of states $\mathcal{X}_{a}\subseteq\mathbb{R}^{n}$ and the safe set $\mathcal{X}_{s}\subseteq\mathbb{R}^{n}$, if $\mathcal{X}_{s}$ is forward invariant and $\mathcal{X}_{s}\subseteq\mathcal{X}_{a}$. We say that a controller $u_{cl}(x)$ guarantees safety if the corresponding closed-loop system is safe.
Assumption 1 (standing assumption)
In the following, the allowable set $\mathcal{X}_{a}$ is assumed compact and regular.
II-B Putinar’s Positivstellensatz
Before stating Putinar’s P-satz, we need to introduce two preliminary definitions.
Definition 1
A quadratic module generated by polynomials $f_{2},...,f_{n}\in R[x]$ is defined by
$\textbf{M}(f_{2},...,f_{n}):=\left\{\gamma_{0}+\sum_{i=2}^{n}\gamma_{i}f_{i}\mid\gamma_{0},...,\gamma_{n}\in\Sigma[x]\right\}.$ (4)
Definition 2
A quadratic module $\textbf{M}(f_{2},...,f_{n})$ is said to be Archimedean when there exists a polynomial $f\in\textbf{M}(f_{2},...,f_{n})$ such that $\{x\in\mathbb{R}^{n}\mid f\geq 0\}$ is a compact set.
The following theorem states Putinar’s P-satz (cf. [17, Theorem 3.20]).
Theorem 1
Given polynomials $f_{1},...,f_{n}\in R[x]$ such that the quadratic module $\textbf{M}(f_{2},...,f_{n})$ is Archimedean, and
$\begin{split}\Bigl{\{}x\mid&f_{1}(x)\leq 0,f_{2}(x)\geq 0,...,f_{n}(x)\geq 0\Bigr{\}}=\emptyset,\end{split}$ (5)
then
$f_{1}\in\textbf{M}(f_{2},...,f_{n}).$ (6)
In words, if the set in 5 is empty, Putinar’s P-satz ensures that the polynomials $\gamma_{0},\gamma_{2},...,\gamma_{n}$ defined in 4 exist, albeit without specifying their degree. To find these polynomials computationally requires to iteratively increase the degree of the polynomials until a solution can be found for 6. An increase in the degree of the polynomials will, however, deteriorate computation time. For many practical examples of interest, however, low-degree polynomials suffice.
Condition 6 can also be expressed more explicitly through the condition that there exists $\gamma_{1},\gamma_{2},...,\gamma_{n}\in\Sigma[x]$ such that
$\gamma_{1}f_{1}-\gamma_{2}f_{2}-...-\gamma_{n}f_{n}\in\Sigma[x].$ (7)
Assuming the empty-set condition in 5 holds, a solution to 7 involving $\gamma_{1}(x)=1$ is guaranteed by Putinar’s P-satz to exist when the degrees of the polynomials are possibly unbounded. In this case, the introduction of the SOS polynomial $\gamma_{1}(x)$ needlessly increases the set of feasible polynomials $f_{1}(x)$. However, when the degrees of the polynomials are truncated, this assertion may no longer hold true, justifying the utilization of $\gamma_{1}(x)$. For the rest of the paper, we refer to both inclusions, 6 and 7, as an SOS constraint.
III Specifications and Problem statement
III-A Specifications
The concept of a safety filter (see Figure 2) is based on the idea of enhancing the behavior of a previously designed legacy controller $u_{n}(x)$ with safety guarantees. The safety filter adjusts the control action of the legacy controller, ideally in a minimal way, to render the system safe with respect to an allowable set $\mathcal{X}_{a}$, encoding the state constraints, and a safe set $\mathcal{X}_{s}$. The safe set denotes a set of states that is forward invariant for some state feedback controller. Therefore, we require that the safe set is strictly contained within the allowable set, i.e.,
$\mathcal{X}_{s}\subseteq\mathcal{X}_{a}.$ (8)
The safe set is preferably chosen large in volume to closely approximate the allowable set.
Figure 2: The safety filter $u_{s}(u_{n},x)$ adjusts the legacy control action $u_{n}(x)$ – if necessary – to guarantee safe operation.
The standard characterization of a safety filter does not specify the extent to which it is allowed to alter the legacy control action. This can be problematic, as the safety filter might inadvertently undermine the desired behavior of the legacy controller. In contrast, our advanced safety filter $u_{s}(x)$ specifies a distinct region $\mathcal{X}_{n}$, where the safety filter remains inactive, that is, for all $x\in\mathcal{X}_{n}$:
$u_{s}(x)=u_{n}(x).$ (9)
This nominal region $\mathcal{X}_{n}$ is required to be forward invariant under the legacy control $u_{n}(x)$, and to be strictly contained within the safe set, i.e.,
$\mathcal{X}_{n}\subsetneq\mathcal{X}_{s}.$ (10)
For a legacy controller that stabilizes the closed-loop system to an attractor within $\mathcal{X}_{a}$, a forward invariant set does always exist.
Moreover, the advanced safety filter should preserve such a stability guarantee across the entire safe set. This is already accomplished within the nominal region $\mathcal{X}_{n}$ due to its forward invariance property. Hence, we only need to additionally ensure a finite-time convergence to $\mathcal{X}_{n}$ for all trajectories originating within the transitional region, defined as:
$\mathcal{X}_{t}:=\mathcal{X}_{s}\setminus\text{int}(\mathcal{X}_{n}),$ (11)
where $\text{int}(\mathcal{X}_{n})$ denotes the interior of a set $\mathcal{X}_{n}$.
Remark 1
The volume of $\mathcal{X}_{t}$ must be balanced with the volume of $\mathcal{X}_{n}$. A larger $\mathcal{X}_{t}$ ensures a smooth behavior of the safety filter when returning the system to $\mathcal{X}_{n}$, while a larger $\mathcal{X}_{n}$ provides a wider nominal region where the safety filter does not modify the legacy control action.
III-B CBF and CLF conditions formalizing the specifications
Figure 3: The advanced safety filter is characterized by the following specifications: (a) the safe set $\mathcal{X}_{s}$ is forward invariant as encoded in 15, (b) it ensures finite-time convergence towards the nominal region $\mathcal{X}_{n}$ as encoded in 17, and (c) the nominal region is forward invariant under the legacy controller $u_{n}(x)$ as encoded in 19 assuming $d(x)=0$.
We employ control Lyapunov and barrier functions to encode the specification from the previous subsection by introducing energy-like scalar functions. In the process, we define the safe set as the intersection of zero-sublevel sets of multiple continuously differentiable functions to closely fit the algebraic structure of the allowable set
$\mathcal{X}_{a}:=\Bigl{\{}x\mid w_{i}(x)\leq 0\text{ }\forall i\in\mathcal{I}_{t}\Bigr{\}}=\bigcap_{i\in\mathcal{I}_{t}}\mathcal{X}_{a,i},$ (12)
where $w_{i}\in R[x]$, $i\in\mathcal{I}_{t}$, are scalar polynomials, and their zero-sublevel sets $\mathcal{X}_{a,i}:=\{x\mid w_{i}(x)\leq 0\}$ are assumed to be compact. This compactness condition is important for applying Putinar’s P-satz in the following section of the paper, and it is not restrictive for the applications we have in mind.
The forward invariance property of the safe set can be asserted by a scalar function $B(x)$ defined as the point-wise maximum of a set of continuously differentiable functions $B_{i}:\mathbb{R}^{n}\to\mathbb{R}$ indexed by $\mathcal{I}_{t}$:
$B(x):=\underset{i\in\mathcal{I}_{t}}{\text{max}}\,B_{i}(x),$ (13)
such that the safe set is given by the intersection of the individual zero-sublevel sets:
$\mathcal{X}_{s}:=\left\{x\mid B(x)\leq 0\right\}=\bigcap_{i\in\mathcal{I}_{t}}\mathcal{X}_{s,i},$
where $\mathcal{X}_{s,i}:=\left\{x\mid B_{i}(x)\leq 0\right\}$. Taking the intersection of multiple functions $B_{i}(x)$ allows for greater flexibility to increase the volume of the safe set compared to using the zero-sublevel set of a single CBF function. This is because the zero-sublevel sets of each $B_{i}(x)$ can extend beyond the allowable set, provided their intersection remains within the allowable set. In contrast, the zero-sublevel set of a single CBF must be fully contained within the allowable set. Hence, we encode the containment condition in 8 as:
$\mathcal{X}_{s,i}\subseteq\mathcal{X}_{a,i},$ (14)
for every $i\in\mathcal{I}_{t}$. This can be seen from $\bigcap_{i\in\mathcal{I}_{t}}\mathcal{X}_{s,i}\subseteq\bigcap_{i\in\mathcal{I}_{t}}\mathcal{X}_{a,i}$ implying $\mathcal{X}_{s}\subseteq\mathcal{X}_{a}$. Here, we assume that the CBF $B(x)$ is defined using the same number of polynomials $B_{i}(x)$ as polynomials $w_{i}(x)$ that define the allowable set. To reduce the number of decision variables in $B(x)$, we can equate certain polynomials $B_{i}(x)$ in the resulting SOS problem. This reduction results in a safe set comprising fewer distinct subsets $\mathcal{X}_{s,i}$ compared to the subsets $\mathcal{X}_{a,i}$ that define the allowable set.
Assumption 2
We assume the safe set $\mathcal{X}_{s}$ to be regular in the sense that $\forall x\in\mathcal{X}_{s},\exists z\in\mathbb{R}^{n},\forall i\in\mathcal{I}_{t}$ [18, Definition 4.9]:
$B_{i}(x)+\nabla B_{i}(x)^{T}z<0$
The regularity assumption in Assumption 2 is a mild one, and it generically holds true, if $B_{i}(x)$ is chosen as a polynomial. It implies that the interior of the safe set is given by $\text{int}(\mathcal{X}_{s})=\left\{x\mid B_{i}(x)<0\quad i\in\mathcal{I}_{t}\right\}$. Examples of polynomials that violate the regularity assumption can be found in [18, Chapter 4.2].
To ensure forward invariance of $\mathcal{X}_{s}$, as shown in Figure 3(a), $B(x)$ needs to satisfy Nagumo’s theorem, that is, $\forall x\in\partial\mathcal{X}_{s},\exists u\in\mathbb{R}^{m}$ (cf. [18, Nagumo’s Theorem 4.7]):
$f(x)+G(x)u\in\mathcal{T}_{\mathcal{X}_{s}}(x),$ (15)
where, under Assumption 2, the tangent cone is defined by
$\mathcal{T}_{\mathcal{X}_{s}}(x):=\{z\mid\nabla B_{i}(x)^{T}z\leq 0\quad i\in\text{Act}(x)\},$ (16)
and $\text{Act}(x):=\left\{i\in\mathcal{I}_{t}\mid B_{i}(x)=0\right\}$ denotes the set of active constraints. The function $B(x)$ in 13 satisfying 15 and 16 will henceforth be referred to as a Control Barrier Function (CBF).
Compared to the standard safety filter, our approach also require finite-time convergence towards $\mathcal{X}_{n}$, as outlined in the specifications. This is asserted by a Control Lyapunov Function (CLF) defined by a continuously differentiable function $V:\mathbb{R}^{n}\to\mathbb{R}$. This function needs to satisfy a Control Lyapunov-like condition in the transitional region $\mathcal{X}_{t}$, as shown in Figure 3(b). Namely, $\forall x\in\mathcal{X}_{t},\exists u\in\mathbb{R}^{m}$:
$\nabla V(x)^{T}\left(f(x)+G(x)u\right)+d(x)\leq 0$ (17)
with dissipation rate
$d(x):=\lambda(x)\left(V(x)+V_{0}\right)>0,$ (18)
where $\lambda:\mathbb{R}^{n}\to\mathbb{R}_{>0}$ is a Lipschitz continuous function, and $V_{0}\leq\text{min}_{x}\left(V(x)\right)$ is a constant value, ensuring $d(x)$ is strictly positive (and thus $V(x)$ strictly decaying) in the transitional region $\mathcal{X}_{t}$. By utilizing a strictly positive function $d(x)$, the CLF condition can be expressed as a non-strict inequality condition similar to the CBF condition – a crucial distinction for the subsequent arguments presented in the next section. We define the nominal region as the zero sublevel set of $V(x)$:
$\mathcal{X}_{n}:=\left\{x\mid V(x)\leq 0\right\}.$
Consequently, the transitional region is given by:
$\mathcal{X}_{t}:=\left\{x\mid B(x)\leq 0\leq V(x)\right\}.$
The next lemma establishes the finite-time convergence property associated with the CLF $V(x)$. Its proof can be found in the appendix.
Lemma 1
Given a Lipschitz-continuous controller $u_{s}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ that satisfies condition 17 with $u:=u_{s}(x)$ and renders the safe set $\mathcal{X}_{s}$ forward invariant, then for all trajectories of system 3 starting in the safe set $\mathcal{X}_{s}$ there exists some finite $T\geq 0$ such that $x(t)\in\mathcal{X}_{n}$ for all $t\geq T$.
Additionally, to make the nominal region forward invariant w.r.t. the legacy controller $u_{n}(x)$, as depicted in Figure 3(c), we propose the condition that $\forall x\in\partial\mathcal{X}_{n}$:
$\nabla V(x)^{T}\left(f(x)+G(x)u_{n}(x)\right)+d(x)\leq 0.$ (19)
Notice that forward invariance also follows with $d(x)=0$. By adopting a slightly stricter forward invariance condition in 19 involving the strictly positive dissipation rate $d(x)$, we ensure the existence of the safety filter implementation as specified in Section V.
Remark 2
As the volume of the transitional region $\mathcal{X}_{t}$ approaches zero, the proposed safety filter transitions into a safety filter based on a single CBF $V(x)=B_{1}(x)=\ldots=B_{t}(x)$, where the CBF and CLF condition 15 and 17 are automatically satisfied by the forward invariance condition 19 of the legacy controller. Therefore, feasible CBF and CLF candidates can be found (in the limit) if a forward invariant set for the system, governed by the legacy controller, exists and is fully contained within the allowable set.
III-C Problem statement
The challenge of deriving the advanced safety filter can be divided into two subproblems. The first subproblem resolves around finding $B(x)$ and $V(x)$ that fulfill the CBF and CLF condition outlined in the preceding section. This involves constructing a rational controller $u_{\text{SOS}}(x)$ to prove the compatibility of the CLF and CBF conditions in 15 and 17. Furthermore, we necessitate the assurance of the forward invariance condition of the legacy controller as defined in 19.
Problem 1
Given a dynamical system 3, an allowable set of states $\mathcal{X}_{a}$ and a legacy controller $u_{n}(x)\in\mathcal{U}$, find scalar functions $B(x):\mathbb{R}^{n}\to\mathbb{R}$ and $V(x):\mathbb{R}^{n}\to\mathbb{R}$ such that:
  • (set containment) the containment conditions in 8 and 10 are satisfied,
  • (finite-time convergence) the CLF condition in 17 is satisfied,
  • (forward-invariance of $\mathcal{X}_{s}$) CBF and CLF conditions in 15 and 17 are satisfied on $\partial\mathcal{X}_{s}$ for the same $u$,
  • (forward-invariance of $\mathcal{X}_{n}$) legacy control condition in 19 is satisfied,
  • (objective) CBF $B(x)$ maximizes the volume of the safe set, denoted by $\text{vol}(\mathcal{X}_{s})$.
For the second subproblem, a safety filter $u_{s}(x)$ is constructed from $B(x)$ and $V(x)$, which is guaranteed to be feasible and Lipschitz-continuous, entailing a smooth transitioning from the safety filter operation to the legacy control operation.
Problem 2
Given a dynamical system 3, a Lipschitz-continuous legacy controller $u_{n}(x)$, a CBF $B(x)$ and CLF $V(x)$, as specified in Problem 1, find state-feedback control law $u_{s}(x)$ that is feasible for all $x\in\mathbb{R}^{n}$, satisfies the CBF and CLF conditions 15 and 17 with $u:=u_{s}(x)$, coincides with $u_{n}(x)$ on $\mathcal{X}_{n}$ (cf. 9), and is Lipschitz-continuous.
In Section IV, we delve into a solution for Problem 1, while a solution for Problem 2 is derived in Section V.
Figure 4: Problem 1 involves searching for compatible CBF $B(x)$ and CLF $V(x)$ using SOS tools. Based on $B(x)$ and $V(x)$, the linear constraints of the QCQP given by $C(x)$, $b(x)$ and $r(x)$ are determined.
IV Constructing CBF and CLF using SOS optimization
IV-A SOS formulation of CBF and CLF conditions
The objective of this subsection is to derive sufficient SOS conditions that encode the CBF condition on $\mathcal{X}_{s}$ and $\mathcal{X}_{n}$, as specified in 15 and 19, and the CLF condition, as specified in 17. We further demonstrate that these SOS conditions are not overly conservative and pose no practical concerns.
The CBF condition in 15 involves the set of active constraints $\text{Act}(x)$, a construction that cannot be directly expressed as an SOS constraint. To address this issue, we first consider the polynomials $B_{i}(x)$, $i\in\mathcal{I}_{t}$, defining the CBF $B(x)$ in 13, as independent CBFs in their own right. By ensuring compatibility among these CBFs $B_{i}(x)$ as proposed in [8], we demonstrate sufficiency of the resulting SOS constraints with respect to CBF condition in 15 (involving the desired CBF $B(x)$). The CBF condition for each $B_{i}(x)$, $i\in\mathcal{I}_{t}$, is expressed as $\forall x\in\{x\in\mathcal{X}_{s}\mid B_{i}(x)=0\},\exists u\in\mathbb{R}^{m}$:
$\nabla B_{i}(x)^{T}\left(f(x)+G(x)u\right)\leq 0.$ (20)
Another challenge in converting conditions 20 and 17 to SOS constraints is the presence of the existential quantifier $\exists u\in\mathbb{R}^{m}$. However, following the approach in [11], we can reformulate these conditions without the need for an existential quantifier. The CBF condition 20 can be expressed as $\forall x\in\{x\in\mathcal{X}_{s}\mid B_{i}(x)=0,\nabla B_{i}(x)^{T}G(x)=0\}$:
$\nabla B_{i}(x)^{T}f(x)\leq 0.$ (21)
Similarly, the CLF condition in 17 can be formulated as $\forall x\in\{x\in\mathcal{X}_{t}\mid\nabla V(x)^{T}G(x)=0\}$:
$\nabla V(x)^{T}f(x)+d(x)\leq 0.$ (22)
In order to render the problem tractable, we restrict our attention to polynomial candidates $B_{i}\in R[x]$, $i\in\mathcal{I}_{t}$, and polynomial candidate $V\in R[x]$. We then apply Putinar’s P-satz, as stated in Theorem 1, on a slightly more stringent conditions compared to 19, 21 and 22, where the inequality is replaced by a strict inequality. These more stringent conditions are summarized as the following empty-set conditions, $i\in\mathcal{I}_{t}$:
$\begin{split}\{x\mid&-\nabla B_{i}(x)^{T}f(x)\leq 0,\nabla B_{i}(x)^{T}G(x)=0,\\ &B_{i}(x)=0,-B_{\mathcal{I}_{t}}(x)\geq 0\}=\emptyset\end{split}$ (23a)
$\begin{split}\{x\mid&-\nabla V(x)^{T}f(x)-d(x)\leq 0,\nabla V(x)^{T}G(x)=0,\\ &V(x)\geq 0,-B_{\mathcal{I}_{t}}(x)\geq 0\}=\emptyset\end{split}$ (23b)
$\begin{split}\{x\mid&\nabla V(x)^{T}\left(f(x)+G(x)u_{n}(x)\right)+d(x)\leq 0,\\ &V(x)=0\}=\emptyset,\end{split}$ (23c)
where $B_{\mathcal{I}_{t}}(x)=\begin{bmatrix}B_{1}(x)&\ldots&B_{t}(x)\end{bmatrix}^{T}$ is a vector of CBFs. The resulting SOS constraints are summarized below. A detailed derivation of the SOS constraint 24 from the empty-set condition 23 is given in the supplementary material in [19].
$\begin{split}-\nabla B_{i}^{T}\left(s_{i}f+Gp_{i}\right)&\in\textbf{M}(B_{i},-B_{\mathcal{I}_{t}})\end{split}$ (24a)
$\begin{split}-\nabla V^{T}\left(s_{0}f+Gp_{0}\right)-s_{0}d&\in\textbf{M}(V,-B_{\mathcal{I}_{t}})\end{split}$ (24b)
$\begin{split}-\nabla V^{T}\left(f+Gu_{n}\right)-d&\in\textbf{M}(V,-V),\end{split}$ (24c)
for some $p_{i}\in(R[x])^{m}$ and $s_{i}\in\Sigma[x]$, $i\in\{0\}\cup\mathcal{I}_{t}$. For clarity, we omit the brackets and variables in the function notation. To achieve the simultaneous satisfaction of CBF and CLF conditions 15 and 17 with a single input $u:=p(x)/s(x)$ on $\partial\mathcal{X}_{s}$, as specified in Problem 1, we equate the polynomials $p_{i}(x)$ and $s_{i}(x)$, $i\in\{0\}\cup\mathcal{I}_{t}$. The resulting SOS constraints are given as:
$\begin{split}-\nabla B_{i}^{T}\left(sf+Gp\right)&\in\textbf{M}(B_{i},-B_{\mathcal{I}_{t}})\end{split}$ (25a)
$\begin{split}-\nabla V^{T}\left(sf+Gp\right)-sd&\in\textbf{M}(V,-B_{\mathcal{I}_{t}})\end{split}$ (25b)
$\begin{split}-\nabla V^{T}\left(f+Gu_{n}\right)-d&\in\textbf{M}(V,-V),\end{split}$ (25c)
for $i\in\mathcal{I}_{t}$, and some $p\in(R[x])^{m}$, and $s\in\Sigma[x]$.
Remark 3
The polynomials $p(x)$ and $s(x)$ induce (after dividing 25a and 25b by $s(x)$) a rational controller $u_{\text{SOS}}(x):=p(x)/s(x)$, employed to prove compatibility of the CBFs $B_{i}(x)$ and CLF $V(x)$.
The proof of Lemma 2 can be found in the appendix.
Lemma 2
If $s(x)>0$ and Assumption 2 is satisfied, then the SOS constraints in 25 are sufficient conditions for the CBF and CLF property in 15, 17 and 19 to hold simultaneously with $u:=p(x)/s(x)$.
The SOS constraints in 25 imply the inequalities since $p\in\Sigma[x]$ implies $p(x)\geq 0$:
$\begin{split}\nabla B_{i}^{T}\left(sf+Gp\right)+\gamma_{0,i}B_{i}-\gamma_{1,i}^{T}B_{\mathcal{I}_{t}}&\leq 0\end{split}$ (26a)
$\begin{split}\nabla V^{T}\left(sf+Gp\right)+sd+\gamma_{0,0}V-\gamma_{1,0}^{T}B_{\mathcal{I}_{t}}&\leq 0\end{split}$ (26b)
$\begin{split}\nabla V^{T}\left(f+Gu_{n}\right)+d+\gamma_{n}V&\leq 0,\end{split}$ (26c)
for some $\gamma_{0,i},\gamma_{n}\in R[x]$, $p\in(R[x])^{m}$, $s,\gamma_{0,0}\in\Sigma[x]$, and $\gamma_{1,i},\gamma_{1,0}\in(\Sigma[x])^{t}$, $i\in\mathcal{I}_{t}$.
Lemma 2 only provides sufficient conditions. However, adding an arbitrary small perturbation $\epsilon>0$ to the SOS constraints in 25 results in necessary conditions when restricting to rational controllers $u_{\text{SOS}}(x)=p(x)/s(x)$. This is because 15, 17 and 19 with $u:=u_{\text{SOS}}(x)$ imply strict inequalities involving $\epsilon$ as follows: $p(x)\leq 0$ implies $p(x)-\epsilon<0$. These strict inequalities imply the SOS constraints 25 perturbed by $\epsilon$ according to Theorem 1. Hence, for all practical purposes, the solution set of the SOS constraints in 25 encompasses the desired set of CBFs and CLFs.
IV-B SOS formulation of input constraints
In the following, we consider two types of inputs constraints: linear and quadratic input constraints. Consider the linear input constraint set
$\mathcal{U}=\{u\mid Au\leq b\},$ (27)
where $A\in\mathbb{R}^{m\times l}$ and $b\in\mathbb{R}^{l}$. These linear input constraints can be encoded by the SOS constraints:
$\begin{split}-a_{i}p(x)+b_{i}s(x)\in\textbf{M}(-B_{\mathcal{I}_{t}}),\end{split}$ (28)
where $a_{i}$ is the $i$th row of $A$, $i\in\mathcal{I}_{l}$, and $p(x)$ and $s(x)$ are as in 25. Next, we consider the quadratic input constraint
$\mathcal{U}=\{u\mid(u-u_{0})^{T}Q_{u}(u-u_{0})\leq u_{\text{max}}^{2}\},$ (29)
for some $Q_{u}\succeq 0$ and $u_{0}\in\mathbb{R}^{m}$. The quadratic input constraint 29 can be encoded using the SOS constraint involving variables $x\in\mathbb{R}^{n}$ and $v\in\mathbb{R}^{m}$:
$\begin{split}&-v^{T}Q_{u}\left(p(x)-s(x)u_{0}\right)+s(x)u_{\text{max}}^{2}\in\textbf{M}(-f_{u},-B_{\mathcal{I}_{t}}),\end{split}$ (30)
where $f_{u}(x,v):=v^{T}Q_{u}v-u_{\text{max}}^{2}\in R[x,v]$ encodes the input constraint set in 29. The following lemma establishes the implication of the input constraints based on the SOS constraints. The proof is reported in the appendix.
Lemma 3
If $s(x)>0$, a solution $p(x)$ and $s(x)$ to the SOS constraint 28, resp. 30, also satisfies the input constraints 27, resp. 29, with $p(x)/s(x)\in\mathcal{U}$ for all $x\in\mathcal{X}_{s}$.
IV-C Safe set and nominal region requirement
In this section, we express the containment conditions of the allowable set $\mathcal{X}_{a}$, the safe set $\mathcal{X}_{s}$ and the nominal region $\mathcal{X}_{n}$, as defined in 14 and 10 (cf. Figure 1), in terms of SOS constraints. A slightly more stringent condition compared to 14, can be expressed as the empty-set condition for all $i\in\mathcal{I}_{t}$:
$\{x\mid B_{i}(x)\leq 0,w_{i}(x)\geq 0\}=\emptyset.$ (31)
Using Putinar’s P-satz in Theorem 1, we translate 31 into the SOS constraints for all $i\in\mathcal{I}_{t}$:
$B_{i}\in\textbf{M}(w_{i}).$ (32)
Similarly, the containment condition of $\mathcal{X}_{n}$ inside $\mathcal{X}_{s}$ in 10, can be expressed as the empty-set condition for all $i\in\mathcal{I}_{t}$:
$\{x\mid V(x)\leq 0,B_{i}(x)\geq 0\}=\emptyset.$ (33)
Using Putinar’s P-satz, we translate 33 into the SOS constraints for all $i\in\mathcal{I}_{t}$:
$-B_{i}\in\textbf{M}(V).$ (34)
IV-D Resulting SOS problem
The solution to Problem 1, i.e., finding suitable functions $B_{i}(x)$, $i\in\mathcal{I}_{t}$, and $V(x)$ can be obtained by solving the following SOS optimization problem:
[…]
where 25 encode the CBF and CLF condition, 28 and 30 encode linear and quadratic input constraints, 32 and 34 encode the set containment properties, and a constant $\epsilon>0$ assures that $s(x)>0$.
IV-E Alternating Algorithm
The solution of the SOS problem 39 can be found by iteratively alternate between searching over one set of decision variables while keeping the others fixed, see [11, 8]. In this process, either constraint margins are minimized or the volume of $\mathcal{X}_{s}$ is maximized. The volume is approximated by the summation of traces of $Q_{B_{i}}$ in $B_{i}(x)=Z_{B,i}(x)^{T}Q_{B,i}Z_{B,i}(x)$, similar to the approach in [7, 20]. Such an alternating algorithm requires a feasible initialization of the decision variables that are kept fixed in the first iteration. This is achieved through an initialization procedure that starts with an initially empty region of interest and expands it until it encompasses the allowable set $\mathcal{X}_{a}$. For more details, we refer the reader to Section IX.
V Safety Filter Implementation
The objective of this section is to implement the advanced safety filter as specified in Problem 2. Given a linear and quadratic input constraints set $\mathcal{U}$, we propose the following realization of the Quadratically Constrained Quadratic Program (QCQP) in closed loop for every state $x\in\mathbb{R}^{n}$ along the trajectory:
$\begin{array}[]{llll}u_{s}(x):=&\underset{u\in\mathcal{U}}{\mbox{min}}&\|u_{n}(x)-u\|^{2}\\ &\mbox{s.t.}&C(x)u+b(x)\leq r(x),\\ \end{array}$ (42)
where
$C(x)$ $:=\begin{bmatrix}\nabla V(x)^{T}G(x)\\ \nabla B_{1}(x)^{T}G(x)\\ \vdots\\ \nabla B_{t}(x)^{T}G(x)\\ \end{bmatrix}$ (43a)
$b(x)$ $:=\begin{bmatrix}\nabla V(x)^{T}f(x)+d(x)\\ \nabla B_{1}(x)^{T}f(x)\\ \vdots\\ \nabla B_{t}(x)^{T}f(x)\end{bmatrix}$ (43b)
encode the CLF and CBF conditions 15 and 17. The smooth functions $r(x)=\begin{bmatrix}r_{1}(x)&\ldots&r_{t}(x)\end{bmatrix}$ are state-dependent slack variables. Within the domain of CLF and CBF conditions 15 and 17, they must satisfy the upper bounds, given by:
$r_{0}(x)\leq 0$ (44)
for all $x\in\mathcal{X}_{t}$, and
$r_{i}(x)\leq 0$ (45)
for all $i\in\mathcal{I}_{t}$ and $x\in\partial\mathcal{X}_{s}$. On the one hand, selecting a larger positive value for $r_{0}(x)$ within $\mathcal{X}_{n}$ (i.e. outside the domain of the CLF condition) increases the allowed dissipation rate violation in terms of larger values of $\dot{V}(x)=\nabla V(x)^{T}\dot{x}$. On the other hand, selecting a larger positive value for $r_{i}(x)$, $i\in\mathcal{I}_{t}$, within $\mathcal{X}_{s}$ (i.e. outside the domain of the CBF condition) increases the allowed safety violation in terms of larger positive values of $\dot{B}(x)=\nabla B(x)^{T}\dot{x}$. This renders the safety filter more passive within the interior of the safe set, and thus leaves more room to satisfy the CLF constraint. However, an overly large value for $r_{i}(x)$, $i\in\mathcal{I}_{t}$, can result in abrupt modifications of the control action when approaching the boundary of the safe set.
Remark 4
In the literature on safety filters, see for example [4], the negative of the allowed dissipation rate and the safety violation are usually specified by class $\mathcal{K}$ functions $\alpha(V(x))$ and $\gamma_{i}(B_{i}(x))$:
$\nabla V(x)^{T}(f(x)+G(x)u)\leq-\alpha(V(x)),$ (46a)
$\nabla B_{i}(x)^{T}(f(x)+G(x)u)\leq-\gamma_{i}(B_{i}(x))\quad i\in\mathcal{I}_{t}.$ (46b)
In this case, the class $\mathcal{K}$ functions $\alpha$ and $\gamma_{i}$ depend on the value of $V(x)$ and $B_{i}(x)$. However, in 43, we consider a broader class of the allowed dissipation rate violation $r_{0}(x)$ and allowed safety violation, $r_{i}(x)$, $i\in\mathcal{I}_{t}$, to ensure a smooth safety filter behavior, while satisfying further requirements presented below.
To ensure perfect tracking of the legacy controller $u_{n}(x)$ within $\mathcal{X}_{n}$ (cf. 9), we further lower bound $r(x)$ as:
$C(x)u_{n}(x)+b(x)\leq r(x)$ (47)
for all $x\in\mathcal{X}_{n}$. Indeed, condition 47 ensures that the cost of the QCQP in 42 can be minimized to zero within $\mathcal{X}_{n}$ with $u=u_{n}(x)$, making $u_{s}(x)$ coincide with $u_{n}(x)\in\mathcal{U}$. To ensure existence of $r(x)$ satisfying lower and upper bounds 44, 45 and 47, we need to show for $r_{0}(x)$ that $\nabla V(x)^{T}\left(f(x)+G(x)u_{n}(x)\right)+d(x)\leq 0$ on the intersection of $\mathcal{X}_{t}$ (cf. 44) and $\mathcal{X}_{n}$ (cf. 47). This is however directly ensured by forward invariance condition 19. Furthermore, we need to show for $r_{i}(x)$, $i\in\mathcal{I}_{t}$, that $\nabla B_{i}(x)^{T}\left(f(x)+G(x)u_{n}(x)\right)\leq 0$ on the intersection of $\partial\mathcal{X}_{s}$ (cf. 45) and $\mathcal{X}_{n}$ (cf. 47). This is however automatically satisfied as the domains do not overlap due to the containment condition 10.
Additionally, the selection of $r(x)$ must be lower-bounded to ensure the feasibility of the QCQP 42 within $\mathcal{X}_{t}$, as 47 already ensures feasibility within $\mathcal{X}_{n}$. Although the exact lower bound is unknown, it can be approximated using a feasible (but not necessarily optimal) solution $u_{\text{sos}}(x)$ from Section IV:
$C(x)u_{\text{sos}}(x)+b(x)\leq r(x),$ (48)
for all $x\in\mathcal{X}_{t}$. To guarantee the existence of such an $r(x)$, note that the lower bound 48 satisfies the upper bound conditions 44 and 45 as a result of the SOS constraints 25a and 25b.
Finally, to establish local Lipschitz-continuity of the proposed safety filter, we need the following assumption.
Assumption 3 (LICQ)
The Linear Independence Constraint Qualification (LICQ) is satisfied for the QCQP in 43 across all states $x$ and inputs $u$, that is, the gradients w.r.t. $u$ of the active inequality constraints in 43 are linearly independent. Equivalently, the rows of $C(x)$ associated with the active inequality constraints are linearly independent.
Remark 5
By extending the QCQP in 43 with a slack variable $\delta\in\mathbb{R}^{t+1}$ as in [21]
$\begin{array}[]{llll}u_{s}(x):=&\underset{(u,\delta)\in\mathcal{U}\times\mathbb{R}^{t+1}}{\mbox{argmin}}&\|u_{n}(x)-u\|^{2}+\lambda\|\delta\|^{2}\\ &\mbox{s.t.}&C(x)^{T}u+b(x)\leq\delta,\\ \end{array}$
we automatically satisfy the LICQ property, as demonstrated in [21]. However, the introduction of these slack variables pose the potential risk of compromising safety and finite-time convergence guarantees. That being said, the LICQ assumption can easily be checked, and it generically holds true, i.e., for almost all polynomial coefficients, the Jacobian of a set of polynomials has full rank.
The local Lipschitz-continuity not only preserves existence and uniqueness of solutions (cf. [15][Theorem 3.1]), but can also avoid the resulting feedback controller from chattering, as explained in [22]. The following theorem establishes local Lipschitz-continuity of the developed QCQP controller. Its proof can be found in the appendix.
Theorem 2
Under Assumption 3, given a smooth $B(x)$ and $V(x)$ as specified in Problem 1, a smooth legacy controller $u_{n}(x)$, and twice differentiable $r_{i}(x)$, $i\in\{0\}\cup\mathcal{I}_{t}$, lower bounded by 48, then $u_{s}(x)$ as defined in 43 is locally Lipschitz-continuous on the safe set $\mathcal{X}_{s}$.
We can find polynomial candidates $r(x)$ based on given CBF $B(x)$ and CLF $V(x)$ by solving an additional SOS problem that encodes the upper and lower bounds 44, 45, 47 and 48.
VI Power Converter Case Study
The proposed safety filter is implemented for a vector field realization of an ac/dc power converter model connected to an infinite bus, as illustrated in Figure 5. A feedback controller that avoids unsafe states is crucial for this application since all electrical variables (e.g. voltages and currents) need to be constrained at all times. For more details about the modelling of power converters, we refer the reader to [23].
Figure 5: The schematic is presented in the (dq) rotating frame, compromising a DC-link, 3-phase bridge, and an ac-line connected to an infinite bus $v_{g}$. Assuming the average-value model for the 3-phase bridge, the input $\bar{u}=\begin{bmatrix}m_{d}&m_{q}\end{bmatrix}^{T}$ determines the converter voltage $v_{c}$ as a function of the dc-link voltage $v_{dc}$. Likewise, the dc-current flowing into the capacitor $c_{dc}$ is given by the input $\bar{u}$ and the ac-line current $i=\begin{bmatrix}i_{d}&i_{q}\end{bmatrix}^{T}$. The states consisting of the dc-link voltage $v_{dc}$ and the line-current $i$ are measured.
The signals are represented in the (dq) rotating frame aligned with the phase $\theta=\omega_{n}t$ of the grid voltage $v_{g}$, where $\omega_{n}=2\pi 50$ is the nominal angular frequency in Hz. The signals are given in per-unit representation. The system’s dynamics are given as follows:
$(1/\omega_{n})c_{dc}\dot{v}_{dc}$ $=-g_{dc}v_{dc}-i_{d}m_{d}-i_{q}m_{q}$ (49a)
$(1/\omega_{n})l\dot{i}_{d}$ $=\omega li_{q}+v_{dc}m_{d}-1$ (49b)
$(1/\omega_{n})l\dot{i}_{q}$ $=-\omega li_{d}+v_{dc}m_{q},$ (49c)
where $l=0.01$ is the inductance in p.u. consisting of transformer and line inductance, $c_{dc}=0.02$ is the dc-link capacitance in p.u., and $g_{dc}:=1/r_{dc}=0.001$ is the admittance of the DC-link capacitor in p.u. The measured states are denoted by $\bar{x}:=\begin{bmatrix}v_{dc}&i_{d}&i_{q}\end{bmatrix}^{T}$, and the input is given by $\bar{u}:=\begin{bmatrix}m_{d}&m_{q}\end{bmatrix}^{T}$, describing the modulation indices. The objective is to stabilize both the DC-link voltage $v_{dc}$ to $1$, and the quadrature current $i_{q}$ to its reference $i_{q,ref}$, which is assumed to be constant. As a consequence, to meet the steady-state specifications, the direct current $i_{d}$ is stabilized to $-g_{dc}$. The allowable set of states must adhere to the following inequality constraints:
$0.2\leq v_{dc}\leq 1.2,\quad i_{d}^{2}+i_{q}^{2}\leq(1.2)^{2}.$ (50)
Furthermore, the quadratic input constraint $\mathcal{U}$ is given by
$m_{d}^{2}+m_{q}^{2}\leq(1.2)^{2}.$ (51)
The system’s equilibrium is characterized by
$x^{*}=\begin{bmatrix}1\\ -g_{dc}\\ i_{q,ref}\end{bmatrix},u^{*}=\begin{bmatrix}1-\omega li_{q,ref}\\ -\omega lg_{dc}\end{bmatrix}.$
We introduce the error coordinates for the states $x=\bar{x}-x^{*}=\begin{bmatrix}\tilde{v}_{dc}&\tilde{i}_{d}&\tilde{i}_{q}\end{bmatrix}^{T}$ and the input variables $u=\bar{u}-u^{*}=\begin{bmatrix}\tilde{m}_{d}&\tilde{m}_{q}\end{bmatrix}^{T}$ such that $\dot{x}=0$ at $x=0$ and $u=0$. The system represented in error coordinates is described by system 2 with
$f(x)=\begin{bmatrix}-0.05\tilde{v}_{dc}-57.9\tilde{i}_{d}+0.00919\tilde{i}_{q}\\ 1710\tilde{v}_{dc}+314\tilde{i}_{q}\\ -0.271\tilde{v}_{dc}-314\tilde{i}_{d}\end{bmatrix}$
$G(x)=\begin{bmatrix}0.05-57.9\tilde{i}_{d}&-57.9\tilde{i}_{q}\\ 1710+1710\tilde{v}_{dc}&0\\ 0&1710+1710\tilde{v}_{dc}\end{bmatrix}.$
We consider a legacy controller that resembles the proportional part of the synchronous rotating frame control structure, as described in [24]:
$u_{n}(x)=\begin{bmatrix}0.1\tilde{v}_{dc}-\tilde{i}_{d}\\ i_{q,ref}-\tilde{i}_{q}\end{bmatrix}.$
In the following subsection, we first consider a single quadrature current reference set to $i_{q,ref}=0$. Then, we introduce a quadrature current reference as a new state to the dynamical system, with a zero time-derivative, i.e., ${\dot{i}}_{q,ref}=0$. For each scenario, we find the CBFs $B_{i}(x)$ and CLF $V(x)$ (cf. Problem 1), and perform simulations that illustrate the additional guarantees provided by the advanced safety filter $u_{s}(x)$ compared with a standard safety filter implementation. We use CVXOPT Python library to solve the SDPs of the alternating algorithm, and to implement the QCQP in 42. This library implements an optimization problem solver based on second-order and positive semidefinite cones. The simulations are performed using the ODE solver from scipy with a sampling time of $1$ $\mu s$. The full code used for simulations and optimization is publicly available online [^1] https://github.com/MichaelSchneeberger/advanced-safety-filter/.
Figure 6: The allowable set $\mathcal{X}_{a}$, the safe set $\mathcal{X}_{s}$, and the nominal region $\mathcal{X}_{n}$ projected to the $\tilde{v}_{dc}$ and $\tilde{i}_{d}$ coordinate are shown for $\tilde{i}_{q}=i_{q,ref}=0$. In addition, the zero-level sets of the polynomials $B_{1}(x)$ and $B_{2}(x)$ are shown in grey. Three scenarios are shown, where (a) the input action is subjected to no constraints, i.e. $u(x(t))\in\mathbb{R}^{m}$, (b) the input action is subjected to a quadratic constraint given by $u(x(t))\in\mathcal{U}:=\{u\mid u^{T}u\leq 1.3\}$, and (c) $i_{q,ref}$ is integrated as a new state of the dynamical system using the same input constraints. The corresponding vector field $\dot{x}=f(x)+G(x)u_{sos}(x)$ (projected to $(\tilde{v}_{dc},\tilde{i}_{d})$ coordinates) is shown in blue, where $u_{sos}(x)=p(x)/s(x)$ is used to ensure compatibility between the CBF and CLF.
VI-A Zero quadrature current reference tracking
In this subsection, we assume a zero quadrature current reference, i.e., $i_{q,ref}=0$. The allowable set $\mathcal{X}_{a}$, as expressed in 12, encoding the state constraints 50, are achieved through the following polynomials:
$w_{1}(x)$ $=((\tilde{v}_{dc}+0.3)/0.5)^{2}+(\tilde{i}_{d}/20)^{2}+(\tilde{i}_{q}/20)^{2}-1$
$w_{2}(x)$ $=(\tilde{v}_{dc}/20)^{2}+({\tilde{i}_{d}}/1.3)^{2}+({\tilde{i}_{q}}/1.3)^{2}-1.$
Figures 6 (a) and (b) illustrates the safe set $\mathcal{X}_{s}$ and the nominal set $\mathcal{X}_{n}$ computed by the algorithm both without and with considering inputs constraints. Table I summarizes both the number of variables associated with the respective SDPs and the number of constraints, derived from the summation of triangular elements within the positive semidefinite matrices. The initialization procedure completed in 10 iterations, where each iteration took around 10 seconds. The main algorithm completed in 20 iterations, where each iteration took around 14 seconds. The degree of the polynomials are given in Table II.
TABLE I: Summary of SDP variables and constraints
Scenario init 1 init 2 main 1 main 2
Fig. 6(a) Variables 475 187 410 475
Constraints 5726 4425 5807 5726
Fig. 6(b) Variables 506 197 441 506
Constraints 6102 4765 6183 6102
Fig. 6(c) Variables 1086 357 981 1086
Constraints 21277 17327 21474 21277
TABLE II: Degree of the polynomials encoding the CLF, CBF, and the rational controller
$V(x)$ $B_{i}(x)$ $p(x)$ $s(x)$
4 4 3 2
The proposed safety filter introduces additional guarantees, which make it more stringent and, as noted in the introduction, more suitable in already established industrial setups. However, due these system requirements, it now represents a new category that prioritizes the preservation of the legacy control, enhancing composability, preventing a direct performance comparison between conventional safety filters and our novel setup. Nonetheless, side-by-side plots in Figure 7 are included that illustrate the additional guarantees compared to a basic safety filter in selected scenarios. The advanced safety filter $u_{s}(x)$ is based on the previously computed CLF $V(x)$ and CBFs $B_{1}(x)$ and $B_{2}(x)$, and is compared to a basic safety filter $u_{\text{SF}}(x)$ using the same CBFs:
$\begin{array}[]{lll}\underset{u\in\mathcal{U}}{\mbox{argmin}}&\|u_{n}(x)-u\|^{2}\\ \mbox{s.t.}&\nabla B_{1}(x)(f(x)+G(x)u)+\alpha\left(B_{1}(x)\right)\leq 0\\ &\nabla B_{2}(x)(f(x)+G(x)u)+\alpha\left(B_{2}(x)\right)\leq 0,\end{array}$ (57)
where $\alpha(B_{i}(x)):=10B_{i}(x)$ is selected.
Figure 7: (Convergence to the nominal region $\mathcal{X}_{n}$): While the advanced safety filter $u_{s}(x)$ essentially overlaps the legacy control $u_{n}$, the basic safety filter $u_{\text{SF}}(x)$ suboptimally perturbs the legacy control action, resulting in a much slower convergence.
Moreover, a side-by-side plots in Figure 8 illustrate the effect on including quadratic input constraints into the design of the safety filter. We consider the integration of quadratic input constraints into the SOS constraint as a contribution distinct from the discussion on the safety filter’s role in preserving the legacy control action. However, for consistency, the simulations were performed using the previous setup, which including an advanced safety filter and the basic safety filter in 57. Since the input constraints are incorporated into the design of the CLF and CBFs, the safety filter $u_{s}(x)$ successfully stays within the state constraints for all $t\geq 0$. This is not true for a basic safety filter $u_{\text{SF}}(x)$, which does not consider the input constraints in its design. In this case, the norm of the input action $\|u_{\text{SF}}(x)\|$ exceeds the input constraint limit of $1.3$ p.u. in the beginning of the simulation (cf. Figure 8). To produce a valid input, the input action $u_{\text{SF}}(x)$ is projected to the set of feasible input set $\mathcal{U}$. As a result, the trajectory deviates from its intended path, leading to a violation of the dc-link voltage constraint.
Figure 8: (Forward invariance of the safe set $\mathcal{X}_{s}$): While the advanced safety filter $u_{s}(x)$, which considers the quadratic input constraints into the design, successfully stays within the state constraints, whereas the legacy controller $u_{n}$ and basic safety filter $u_{\text{SF}}(x)$ violate the state constraints.
VI-B Including current reference tracking
For this subsection, the objective is to stabilize both the DC-link voltage to $1$, and the quadrature current to some (not yet a priori pre-specified) quadrature current reference $i_{q,ref}$. To render our design parametric in $i_{q,ref}$, we introduce a new state representing the quadrature current reference. Consequently, we augment the dynamics from 49 with $\dot{i}_{q,ref}=0$.
Remark 6
Although the reference $i_{q,ref}$ is assumed constant for the design of the CBFs and CLF, step-wise adjustments to $i_{q,ref}$ over time (corresponding to task cycles of the digital processing unit) do not compromise safety, provided the state remains within the safe set $\mathcal{X}_{s}$.
The corresponding coordinates of the states and inputs are given by $x=\begin{bmatrix}\tilde{v}_{dc}&\tilde{i}_{d}&\tilde{i}_{q}&\tilde{i}_{q,ref}\end{bmatrix}^{T}$ and $u:=\begin{bmatrix}\tilde{m}_{d}&\tilde{m}_{q}\end{bmatrix}^{T}$, where $\tilde{i}_{q,ref}=i_{q,ref}$. The allowable set $\mathcal{X}_{a}$ is chosen through the following polynomials:
$w_{1}(x)$ $=((\tilde{v}_{dc}+0.3)/0.5)^{2}+(\tilde{i}_{d}/20)^{2}+(\tilde{i}_{q}/20)^{2}$
$+(\tilde{i}_{q,ref}/20)^{2}-1$
$w_{2}(x)$ $=(\tilde{v}_{dc}/20)^{2}+({\tilde{i}_{d}}/1.3)^{2}+({\tilde{i}_{q}}/1.3)^{2}$
$+(\tilde{i}_{q,ref}/20)^{2}-1.$
Correspondingly, a safe set $\mathcal{X}_{s}$ and the nominal region $\mathcal{X}_{n}$ can be computed for this 4-dimensional space. Figure 6(c) illustrates the safe set $\mathcal{X}_{s}$ and the nominal set $\mathcal{X}_{n}$ for $i_{q}=i_{q,ref}=0$. Table I summarizes the number of variables of the corresponding SDP for each part, as well as the number of constraints. The initialization procedure completed in 51 iterations, where each iteration took around 60 seconds. The main algorithm completed in 20 iterations, where each iteration took around 90 seconds. The degree of the polynomials are summarized in Table II.
The simulation shown in Figure 9 illustrates the reference tracking performance of our safety filter $u_{s}(x)$ compared to the basic safety filter $u_{\text{SF}}(x)$ introduced in the previous subsection. The safety filter, although slightly perturbing the legacy controller, guarantees finite-time convergence to the nominal set defined for a quadrature current reference $i_{q,ref}$ of $1$. However, the basic safety filter $u_{\text{SF}}(x)$ perturbs the legacy control action, leading to a much slower convergence. In general, as we approach the boundary of the safe set by increasing $i_{q,ref}$, the basic safety filter is more likely to intervene the legacy control action.
Figure 9: (Quadrature current reference tracking): The closed-loop simulation demonstrates the finite-time convergence guarantee of the advanced safety filter $u_{s}(x)$ towards the nominal region $\mathcal{X}_{n}$ for non-zero quadrature current reference $i_{q,ref}=1$p.u. Additionally, the advanced safety filter exhibits a faster convergence to the equilibrium compared to the basic safety filter $u_{\text{SF}}(x)$. The latter is more likely to disturb the legacy control action as the quadrature current reference $i_{q,ref}$ approaches the boundary of the safe set, resulting in a slower convergence.
VII Conclusion
We successfully developed and implemented a novel safety filter approach that, to our knowledge, is the first of its kind to leave the existing legacy controller subject to input constraints unaffected. This innovation could lead to greater acceptance of safety filters in industrial environments and real-world scenarios, advancing our understanding of their implementation across a broader range of applications. Our next step is to integrate this safety filter with our industrial partners for the control of an advanced converter topology.
VIII Appendix
VIII-A Proof of Lemma 1
Because the safe set $\mathcal{X}_{s}$ is compact and forward invariant, and the closed-loop dynamical system $\dot{x}=f(x)+G(x)u_{s}(x)$ is Lipschitz-continuous, there exists a unique trajectory $x(t)$ for all $x(0)\in\mathcal{X}_{s}$ and $t>0$. The CLF condition in 17 with $u:=u_{s}(x)$ can be written without dissipation rate as $\forall x\in\partial X_{n}$:
$\nabla V(x)^{T}\left(f(x)+G(x)u_{s}(x)\right)\leq 0.$ (58)
Using the tangent cone $T_{\mathcal{X}_{n}}(x)=\{z\mid\nabla V(x)^{T}z\leq 0\}$, the condition 58 can be formulated as $\forall x\in\partial X_{n}$:
$f(x)+G(x)u_{s}(x)\in\mathcal{T}_{\mathcal{X}_{n}}(x).$
Hence, according to [18, Nagumo’s Theorem 4.7], $\mathcal{X}_{n}$ is forward invariant. This proves that for all $x(0)\in\mathcal{X}_{n}$, we can choose $T=0$. lt remains to show that for all other initial states, i.e., $x(0)\in\mathcal{X}_{t}$, there exists $T\geq 0$ such that $x(T)\in\mathcal{X}_{n}$, or equivalently $V\left(x(T)\right)\leq 0$.
As $\mathcal{X}_{t}\subseteq\mathcal{X}_{s}$ is compact (cf. Assumption 1) and the dissipative rate $d(x)$ defined in 18 is assumed to be strictly positive and Lipschitz-continuous on $\mathcal{X}_{t}$, the following minimum exists: $l=\underset{x\in\mathcal{X}_{t}}{\text{min}}\lambda(x)\left(V(x)+V_{0}\right)>0$. Therefore, the scalar function $V\left(x(t)\right)$ strictly decreases over time for all $x(0)\in\mathcal{X}_{t}$ due to the CLF condition in 17: $\dot{V}=\nabla V(x)\left(f(x)+G(x)u_{s}(x)\right)\leq-l<0$. As $\mathcal{X}_{s}$ is forward invariant, $V\left(x(t)\right)$ is upper bounded by $V\left(x(t)\right)\leq V\left(x(0)\right)-lt$ for all $x(0)\in\mathcal{X}_{t}$. Therefore, there exists some $T\leq V\left(x(0)\right)/l$ such that $V\left(x(T)\right)\leq 0$.
VIII-B Proof of Lemma 2
By dividing 26a and 26b by $s(x)>0$, we obtain
$\begin{split}\nabla B_{i}^{T}\left(f+Gu_{\text{sos}}\right)+\frac{\gamma_{0,i}}{s}B_{i}-\frac{\gamma_{1,i}^{T}}{s}B_{\mathcal{I}_{t}}&\leq 0\end{split}$ (59a)
$\begin{split}\nabla V^{T}\left(f+Gu_{\text{sos}}\right)+d+\frac{\gamma_{0,0}}{s}V-\frac{\gamma_{1,0}^{T}}{s}B_{\mathcal{I}_{t}}&\leq 0,\end{split}$ (59b)
where $u_{\text{sos}}(x)=p(x)/s(x)$.
From 59a, by using the fact that $\gamma_{1,i}(x)\geq 0$, $s(x)>0$, and $B_{i}(x)=0$ and $B_{\mathcal{I}_{t}}(x)\leq 0$ on $\partial\mathcal{X}_{s}$, we infer, for each $i\in\mathcal{I}_{t}$, that for all $x\in\partial\mathcal{X}_{s}$ there exists $u:=u_{\text{sos}}(x)$ such that $\nabla B_{i}^{T}\left(f+Gu_{\text{sos}}\right)\leq 0$ whenever $B_{i}(x)=0$. Or, equivalently, for all $x\in\partial\mathcal{X}_{s}$ there exists $u:=u_{\text{sos}}(x)$ such that $\nabla B_{i}^{T}\left(f+Gu\right)\leq 0$ for all $i\in\text{Act}(x)$, as defined in 16. This proofs the CBF condition in 15 using the definition of the tangent cone under Assumption 2.
From 59b, by using the fact that $\gamma_{0,0}(x)\geq 0$, $\gamma_{1,0}(x)\geq 0$, $s(x)>0$, and $V(x)\geq 0$ and $B_{\mathcal{I}_{t}}(x)\leq 0$ on $\mathcal{X}_{t}$, we deduce that for all $x\in\mathcal{X}_{t}$ there exists $u:=u_{\text{sos}}(x)$ such that $\nabla V^{T}\left(f+Gu\right)+\lambda\bigl{(}V-V_{0}\bigr{)}\leq 0$. This proofs the CLF condition 17.
Finally, from 26c, by using the fact that $V(x)=0$ on $\partial\mathcal{X}_{n}$, we deduce that the inequality $\nabla V^{T}\left(f+Gu_{n}\right)\leq 0$ holds for all $x\in\partial\mathcal{X}_{n}$. This proofs the CBF condition in 19.
VIII-C Proof of Lemma 3
We use the definition of a quadratic module in 4, to rewrite the SOS constraint in 28 as $\forall x\in\mathbb{R}^{n},\exists\gamma_{u}\in(\Sigma[x])^{t}$:
$-a_{i}u_{\text{SOS}}(x)+b_{i}+\frac{\gamma_{u}(x)^{T}}{s(x)}B_{\mathcal{I}_{t}}(x)\geq 0$
for $i\in\mathcal{I}_{l}$, and $u_{\text{SOS}}(x)=p(x)/s(x)$. Since the $\gamma_{u}(x)\geq 0$, $s(x)>0$, and $B_{\mathcal{I}_{t}}(x)\leq 0$ on $\mathcal{X}_{s}$, we deduce, for all $i\in\mathcal{I}_{l}$, $-a_{i}u_{\text{SOS}}(x)+b_{i}\geq 0$ on $\mathcal{X}_{s}$. This concludes the first part of the proof.
By substituting the variable $v$ in 30 with $u-u_{0}$, we can reformulate the SOS constraints by using the definition of a quadratic module as $\forall x\in\mathbb{R}^{n},\exists\gamma_{u,1}\in\Sigma[x,u],\gamma_{u,2}\in(\Sigma[x,u])^{t}$:
$\begin{split}&-\left((u-u_{0})^{T}Q_{u}\left(u_{\text{SOS}}(x)-u_{0}\right)-u_{\text{max}}^{2}\right)\\ &\quad+\frac{\gamma_{u,1}(x,u)}{s(x)}f_{u}(x,u-u_{0})+\frac{\gamma_{u,2}(x,u)^{T}}{s(x)}B_{\mathcal{I}_{t}}\geq 0,\end{split}$ (60)
where $x\in\mathbb{R}^{n}$ and $u\in\mathbb{R}^{m}$ are the variables of the polynomial expressions. Because of its positive semidefiniteness, $Q_{u}$ can be decomposed into $Q_{u}=B_{u}^{T}B_{u}$. Furthermore, since $\gamma_{u,1}(x,u)\geq 0$, $\gamma_{u,2}(x,u)\geq 0$, $s(x)>0$, $B_{\mathcal{I}_{t}}(x)\leq 0$ on $\mathcal{X}_{s}$, and $f_{u}(x,u-u_{0})\leq 0$ on $\mathcal{U}$, condition 60 can be reformulated as $\forall x\in\mathcal{X}_{s},\forall u\in\mathcal{U}$:
$(u-u_{0})^{T}B_{u}^{T}B_{u}(u_{\text{SOS}}(x)-u_{0})\leq u_{\text{max}}^{2}.$ (61)
For any $x\in\{x\in\mathcal{X}_{s}\mid B_{u}(u_{\text{SOS}}(x)-u_{0})=0\}$, we have $(u_{\text{SOS}}(x)-u_{0})^{T}B_{u}^{T}B_{u}(u_{\text{SOS}}(x)-u_{0})=0\leq u_{\text{max}}^{2}$. Hence, $u_{\text{SOS}}(x)\in\mathcal{U}$.
For any other safe state, that is $x\in\{x\in\mathcal{X}_{s}\mid B_{u}(u_{\text{SOS}}(x)-u_{0})\neq 0\}$, there exists a unique $\lambda>0$ such that
$\begin{split}\left(B_{u}\lambda(u_{\text{SOS}}(x)-u_{0})\right)^{T}\left(B_{u}\lambda(u_{\text{SOS}}(x)-u_{0})\right)&=u_{\text{max}}^{2}.\end{split}$ (62)
Additionally, substituting $u$ by $\lambda(u_{\text{SOS}}(x)-u_{0})+u_{0}$ in condition 61 yields the condition that $\forall x\in\mathcal{X}_{s}$:
$\lambda(B_{u}(u_{\text{SOS}}(x)-u_{0}))^{T}(B_{u}(u_{\text{SOS}}(x)-u_{0}))\leq u_{\text{max}}^{2}.$ (63)
By comparing 62 and 63, we conclude that $\lambda$ must be greater than $1$. As a consequence, $(B_{u}(u_{\text{SOS}}(x)-u_{0}))^{T}(B_{u}(u_{\text{SOS}}(x)-u_{0}))\leq u_{\text{max}}^{2}$ holds for all $x\in\mathcal{X}_{s}$. This concludes that $u_{\text{SOS}}(x)\in\mathcal{U}$ for all $x\in\mathcal{X}_{s}$.
VIII-D Proof of Theorem 2
Consider the optimization problem
$\begin{array}[]{ll}\mbox{min}_{u}&f(u,x)\\ \mbox{s.t.}&g(u,x)\leq 0,\end{array}$ (66)
and a point $x^{*}\in\mathbb{R}^{n}$. We say $(u_{s}(x^{*}),\lambda(x^{*}))$ is a regular minimizer if the LICQ is satisfied at $u_{s}(x^{*})$, and $(u_{s}(x^{*}),\lambda(x^{*}))$ satisfies the KKT condition and the strong second order sufficient condition (SSOSC). The point $(u,\lambda)$ satisfies the SSOSC if Hessian matrix of the Lagrangian of 66 is positive definite at this point. According to [25, Theorem 2.3.3], there exists a locally Lipschitz continuous map $u_{s}(x)$ around a neighborhood of $x^{*}$ such that $u_{s}(x)\in\mathbb{R}^{m}$ is a minimizer for the non-linear optimization problem 66, if $f$ and $g$ are twice continuously differentiable in $(u,x)$, and $u_{s}(x^{*})$ is a regular minimizer with multiplier $\lambda(x^{*})$. In what follows, we establish these conditions for the QCQP in 43.
LICQ property under Assumption 3 asserts that for any solution of the QCQP $u_{s}(x^{*})$, there exists a unique $\lambda^{*}$ that satisfies the KKT condition at $(u_{s}(x^{*}),\lambda^{*})$. Next, we show that the SSOSC holds for any choice of $\lambda^{*}$, specifically $\lambda^{*}$ that satisfies the KKT condition. Consider the Lagrangian of the QCQP for either linear or quadratic input constraints:
$\begin{split}L(u,\lambda)=&u^{T}u+2u_{n}(x)u+u_{n}(x)^{T}u_{n}(x)\\ &+\lambda_{1}^{T}(C(x)u+b(x))\\ &+\lambda_{2}^{T}(u^{T}Q_{u}u-u_{max}^{2})+\lambda_{3}^{T}(Au-b)\end{split}$
Then the inequality $v^{T}\nabla_{uu}^{2}L(u,\lambda)v=v^{T}v+\lambda_{2}v^{T}Q_{u}v>0$ for all $v\in\mathbb{R}^{m}$. This proves that any solution of the QCQP $u_{s}(x^{*})$ satisfies the SSOSC with the corresponding $\lambda^{*}$. Furthermore, a solution of the QCQP $u_{s}(x)$ exists for all $x$ due to the lower bound on $r(x)$ in 48, consequently establishing the local Lipschitz continuity of $u_{s}(x)$.

### v3

@@
 This paper presents a novel safety filter framework that ensures both safety and the preservation of the legacy control action within a nominal region.
-This modular design allows the safety filter to be integrated into the control hierarchy without compromising the performance of the existing legacy controller within the nominal region.
-This is accomplished by formulating multiple Control Barrier Functions (CBFs) and Control Lyapunov-like Functions (CLFs) conditions, alongside a forward invariance condition for the legacy controller, as sum-of-squares constraints utilizing Putinar’s Positivstellensatz.
-Additionally, the state-dependent inequality constraints of the resulting Quadratic Program – encoding the CBF and CLF conditions – are designed to remain inactive within the nominal region, ensuring perfect tracking of the legacy control action.
+This modular design allows the safety filter to be integrated into the control hierarchy without compromising the performance of the existing legacy controller during nominal operation.
+For a control-affine system, this is accomplished by formulating multiple Control Barrier Functions (CBFs) and Control Lyapunov-like Functions (CLFs) conditions, alongside a forward invariance condition for the legacy controller, as sum-of-squares constraints.
+Additionally, the state-dependent inequality constraints of the resulting Quadratic Program (QP) – encoding the CBF and CLF conditions – are designed to remain inactive within the nominal region, ensuring preservation of the legacy control action and performance.
 Our safety filter design is also the first to include quadratic input constraints, and does not need an explicit specification of the attractor, as it is implicitly defined by the legacy controller.
-To avoid the chattering effect and guarantee the uniqueness and Lipschitz continuity of solutions, the state-dependent inequality constraints of the Quadratic Program are selected to be sufficiently regular.
+To avoid the chattering effect and guarantee the uniqueness and Lipschitz continuity of solutions, the state-dependent inequality constraints of the QP are selected to be regular.
 Finally, we demonstrate the method in a detailed case study involving the control of a three-phase ac/dc power converter.
 I Introduction
-In various engineering applications, it is of much interest to endow a well-performant controller with safety guarantees by minimally changing its control action.
+In several engineering applications, it is of much interest to endow a well-performing controller with safety guarantees by minimally changing its control action.
 This concept is commonly referred to as a safety filter, and a comprehensive overview of the subject can be found in [21].
 An illustrative example is a finely-tuned control design based on linearization, which must be extended to adhere to hard state constraints in the nonlinear closed-loop system.
-We refer to such a well-performant controller as the legacy controller.
-A popular type of safety filters relies on solving a Quadratic Program (QP) at every state using a Control Barrier Function (CBF).
-By continuously solving the QP, the control action of the legacy controller is smoothly adjusted as the system approaches the boundary of the safe set $\mathcal{X}_{s}$.
-The main challenges in deploying such a safety filter lies in finding a viable CBF candidate and adhering to input constraints.
+We refer to such a controller as the legacy controller.
+A popular type of safety filters for control-affine systems relies on a Control Barrier Function (CBF).
+By continuously solving the associate Quadratic Program (QP), the control action of the legacy controller is smoothly adjusted as the system approaches the boundary of the safe set $\mathcal{X}_{s}$.
+The main challenges in deploying such a safety filter lies in finding viable CBF candidates and adhering to input constraints.
 Even though safety filters based on a QP are well-established, they are frequently not adopted in current industrial setups due to their risk of altering the legacy control behavior during nominal operations.
-In practice, the legacy control parameters are often meticulously tuned and thoroughly evaluated, typically involving extensive hardware-in-the-loop simulations, which require significant time and resources.
-As a result, the safety filter is expected to intervene only when the system approaches state constraints, meaning when it is leaving the nominal operating regime.
+In practice, the legacy control parameters are often meticulously tuned and thoroughly evaluated, which require significant time and resources.
+As a result, preserving the legacy control action within the nominal operating regime is not merely a preference but a strict engineering requirement.
+In such contexts, the role of a safety filter is to act as a last-resort mechanism that intervenes only when the system state approaches critical boundaries.
 Figure 1: The safe set $\mathcal{X}_{s}$, defined as the zero sublevel set of the CBF $B(x)$, specifies the set of states for which safety can be guaranteed.
@@
 Unlike in a reach-avoid problem [11], the nominal region $\mathcal{X}_{n}$ does not represent a predetermined set of target states.
-Instead, it is selected during the design process (often at an earlier stage) to encompass the nominal operating regime of the legacy controller while maintaining a sufficiently large transitional region for a smooth safety filter operation.
-The specifications are inherently more stringent than those of a standard safety filter, but may therefore be deployed in an industrial setup without interference in the original line-up.
-Our safety filter is realized based on solving a Quadratic Constraint Quadratic Program (QCQP) at every state using multiple CBFs and Control Lyapunov-like Functions (CLFs).
-A QCQP instead of a simple QP is employed to encode quadratic input constraints, which are common, e.g., in power electronic applications.
-While CBF conditions (encoded as linear constraints of the QCQP) are used to ensure safety w.r.t.
-$\mathcal{X}_{s}$, we leverage the CLF condition (also encoded as linear constraints of the QCQP) to specify a dissipation rate ensuring a finite-time convergence from the boundary of the safe set to the nominal region $\mathcal{X}_{n}$.
-The finite-time convergence property is not intended to enhance control performance but to ensure a rapid return to the nominal operating regime following an unforeseen event.
-We ensure that the nominal set contains the desired attractor by selecting it forward invariant with respect to the legacy controller.
-For identifying compatible CBF and CLF candidates that adhere to the linear and quadratic input constraints, as well as a solution for the QCQP at every state, we propose a Sum-of-Squares (SOS) optimization approach.
-Preserving the legacy control action is achieved by ensuring that the inequality constraints of the QCQP remain inactive within $\mathcal{X}_{n}$.
-This necessitates taking into account the legacy controller when searching for appropriate CBF and CLF candidates.
-To solve the resulting SOS program, we deploy an alternating algorithm that simultaneously searches for a feasible controller in the class of rational functions of the state.
-The concept of our safety filter resembles the integration of a CLF with a CBF to attain both safety and stabilization.
-Such a unification, although just with a single function, has been initially proposed in [3].
-The integration of both CLF and CBF into an optimization-based controller has since become a well-established concept, see the survey in [4].
+Instead, it is selected during the design process (often at an earlier stage) to encompass the nominal operating regime.
+The specifications are inherently more stringent than those of a standard safety filter, but may therefore be deployed in an industrial setup without disrupting the original control architecture.
+Our safety filter is realized for control-affine systems based on solving a QP at every state, where the linear inequality constraints are derived from multiple CBFs, Control Lyapunov-like Functions (CLFs), and state-dependent slack variables.
+These slack variables provide local flexibility by relaxing the constraint in regions where the CBF and CLF condition are not required to hold.
+A Quadratic Constraint Quadratic Program (QCQP) instead of a simple QP can be employed to encode quadratic input constraints, which are common, e.g., in power electronic applications.
+Despite the restriction to control-affine and polynomial systems, our approach remains relevant for a wide range of systems encountered in practice, especially in domains like power systems and robotics, where control-affine polynomial models are commonly used.
+The design process consists of two stages: first, we identify compatible CBF and CLF candidates using a Sum-of-Squares (SOS) optimization approach; second, we determine the state-dependent slack variables via a separate SOS program, ensuring that the legacy control action is preserved within the nominal region.
+Our approach focuses on computing formal guarantees via SOS optimization, enabling an automated design procedure and eliminating the need for application-specific parameter tuning based on extensive simulations.
+The proposed safety filter concept resembles the integration of a CLF with a CBF to attain both safety and stabilization.
+Such an integration, although just with a single function, has been initially proposed in [17], and it has since become a well-established concept, see the survey in [1].
 Feasibility of such an optimization-based controller can be accomplished by introducing a slack variable that deactivates CLFs and CBFs based on their respective priorities, as shown in [2].
@@
 This restriction allows to reformulate the problem with linearity in its decision variables.
-After our previous work [8] on synthesizing a rational controller to find compatible CLF and multiple CBFs using SOS optimization, and through collaboration with industrial partners, we identified a recurring issue: any modifications to the existing legacy control architecture should not compromise the nominal control performance.
-A prime example for this is the integration of a safety filter.
-In this paper, we propose a novel way of implementing a safety filter without disturbing the legacy controller within the nominal region.
-The key advancements are summarized as follows:
-  • First, we propose a novel safety filter implementation that, by design, ensures finite-time convergence to the nominal region $\mathcal{X}_{n}$ and perfect tracking of the legacy control action within that region.
-  • Second, multiple CBFs are used to the increase flexibility of the resulting safe set.
-Instead of imposing each CBF condition separately, we relax these conditions when the corresponding CBF is not active, similar to the idea in [13].
+After our previous work [18] on synthesizing compatible CLF and multiple CBFs using SOS optimization, we identified a recurring issue: industrial standards often requires compliance with pre-certified control behavior, particularly during nominal operations.
+A prime example is the integration of a safety filter.
+To address this, we proposed a modular and minimal invasive safety filter approach that preserves the legacy control behavior within the nominal region, while still ensuring safety near constraint boundaries.
+To address those issues we extended the framework by including the following contributions:
+  • First, we synthesize compatible CLFs and CBFs based on an SOS problem, augmented by an additional SOS constraint that ensures forward-invariance of the nominal region under the legacy control.
+This constraint is essential, as it enables the construction of the proposed safety filter in a subsequent step.
+  • Second, to increase flexibility in shaping the safe set, we employ multiple CBFs.
+Rather than imposing each CBF condition separately, we relax these conditions when the corresponding CBF is not active, similar to the idea in [8].
 Moreover, the CLF condition can also be relaxed by leveraging the forward invariance property of the CBFs, similar to [20] for autonomous systems.
-  • Third, the desired attractor (e.g., an equilibrium point or limit cycle), which is stabilized by the legacy controller in the nominal region, does not need to be specified explicitly, as it is implicitly specified through the legacy controller as a forward invariance condition.
+  • Third, given feasible CLF and CBFs candidates, we formulate a second SOS problem to compute the state-dependent slack variables that ensure safety while preserving the legacy behavior within the nominal region.
   • Fourth, we introduce a novel approach for encoding linear and quadratic input constraints on the rational controller.
-This ensures feasibility of the resulting QCQP-based controller, even in case of quadratic input constraints – an issue that, to the authors’ knowledge, is unsolved.
-To avoid chattering and guarantee the uniqueness and Lipschitz continuity of solutions, the state-dependent inequality constraints of the QCQP are selected to be sufficiently regular.
-Finally, we tested our framework on a detailed three-phase ac/dc converter system connected to an infinite bus, where we successfully synthesized a safety filter that ensures satisfaction of the state constraints given by the dc-link voltage and line current, while adhering to quadratic input constraints.
-We further tracked an a priori unspecified quadrature current reference of the ac/dc converter by integrating the reference as a state into a slightly modified system model.
+This ensures feasibility of the resulting QP-based controller, even in case of quadratic input constraints – an issue that, to the authors’ knowledge, is unsolved.
+To avoid chattering and guarantee the uniqueness and Lipschitz continuity of solutions, the state-dependent inequality constraints of the QP are selected to be regular (see Assumption 2).
+Finally, we tested our framework on a detailed three-phase ac/dc converter system connected to an infinite bus, where we successfully synthesized a safety filter that ensures satisfaction of the state constraints given by the dc-link voltage and line current, while adhering to input constraints.
 The remainder of the paper is structured as follows: In Section II, the preliminaries and notation are discussed.
@@
 II-A Notation
-The shorthand $\mathcal{I}_{t}:=\{1,2,...,t\}$ designates a range of natural numbers from 1 to $t$.
+The shorthand $\mathcal{I}_{n}:=\{1,2,...,n\}$ designates a range of natural numbers from $1$ to $n$.
 $R[x]$ refers to the set of scalar polynomials in variables $x\in\mathbb{R}^{n}$, and $\Sigma[x]$ refers to the set of scalar SOS polynomials in variables $x\in\mathbb{R}^{n}$.
 If $p(x)\in\Sigma[x]$, it can be expanded to
-$p(x)=Z(x)^{T}QZ(x),$ (1)
+$p(x)=Z(x)^{\top}QZ(x),$ (1)
 where $Z(x)$ is a vector of monomials in $x$, and $Q$ is a square positive semidefinite matrix.
+Definition 1
+A quadratic module generated by polynomials $f_{2},...,f_{n}\in R[x]$ is defined by
+$\textbf{M}(f_{2},...,f_{n}):=\left\{\gamma_{0}+\sum_{i=2}^{n}\gamma_{i}f_{i}\mid\gamma_{0},...,\gamma_{n}\in\Sigma[x]\right\}.$ (2)
 Throughout the paper, we consider a polynomial control system, affine in the control action $u\in\mathbb{R}^{m}$, and described as
-$\dot{x}=F_{c}(x,u)=f(x)+G(x)u,$ (2)
+$\dot{x}=f(x)+G(x)u,$ (3)
 where $x\in\mathbb{R}^{n}$ is the state, $f(x)\in\left(R[x]\right)^{n}$ is a polynomial vector field, and $G(x)\in\left(R[x]\right)^{n\times m}$ is a polynomial matrix.
-System 2 with a polynomial state feedback control policy $u_{cl}(x)\in\left(R[x]\right)^{m}$ results in a closed-loop system of the form
+System 3 with a polynomial state feedback control policy $u_{cl}(x)\in\left(R[x]\right)^{m}$ results in a closed-loop system of the form
 $\dot{x}=F_{cl}(x)=f(x)+G(x)u_{cl}(x).$ (4)
@@
 A subset $\mathcal{X}\subseteq\mathbb{R}^{n}$ is called forward invariant (cf.
-[15, Theorem 4.4]) with respect to system 3 if for every $x(0)\in\mathcal{X}$, $x(t)\in\mathcal{X}$ for all $t\geq 0$.
-A system 3 is called safe (cf.
+[10, Theorem 4.4]) with respect to system 4 if for every $x(0)\in\mathcal{X}$, $x(t)\in\mathcal{X}$ for all $t\geq 0$.
+A system 4 is called safe (cf.
 [24]) w.r.t.
 an allowable set of states $\mathcal{X}_{a}\subseteq\mathbb{R}^{n}$ and the safe set $\mathcal{X}_{s}\subseteq\mathbb{R}^{n}$, if $\mathcal{X}_{s}$ is forward invariant and $\mathcal{X}_{s}\subseteq\mathcal{X}_{a}$.
+In the following, the allowable set $\mathcal{X}_{a}$ is assumed compact and regular.
 We say that a controller $u_{cl}(x)$ guarantees safety if the corresponding closed-loop system is safe.
@@
 A quadratic module $\textbf{M}(f_{2},...,f_{n})$ is said to be Archimedean when there exists a polynomial $f\in\textbf{M}(f_{2},...,f_{n})$ such that $\{x\in\mathbb{R}^{n}\mid f\geq 0\}$ is a compact set.
-The following theorem states Putinar’s P-satz (cf. [17, Theorem 3.20]).
+The following theorem states Putinar’s P-satz (cf. [13, Theorem 3.20]).
 Theorem 1
@@
 $f_{1}\in\textbf{M}(f_{2},...,f_{n}).$ (7)
-In words, if the set in 5 is empty, Putinar’s P-satz ensures that the polynomials $\gamma_{0},\gamma_{2},...,\gamma_{n}$ defined in 4 exist, albeit without specifying their degree.
-To find these polynomials computationally requires to iteratively increase the degree of the polynomials until a solution can be found for 6.
+In words, if the set in 6 is empty, Putinar’s P-satz ensures that the polynomials $\gamma_{0},\gamma_{2},...,\gamma_{n}$ defined in 5 exist, albeit without specifying their degree.
+To find these polynomials computationally requires to iteratively increase the degree of the polynomials until a solution can be found for 7.
 An increase in the degree of the polynomials will, however, deteriorate computation time.
 For many practical examples of interest, however, low-degree polynomials suffice.
-Condition 6 can also be expressed more explicitly through the condition that there exists $\gamma_{1},\gamma_{2},...,\gamma_{n}\in\Sigma[x]$ such that
+Condition 7 can also be expressed more explicitly through the condition that there exists $\gamma_{1},\gamma_{2},...,\gamma_{n}\in\Sigma[x]$ such that
 $\gamma_{1}f_{1}-\gamma_{2}f_{2}-...-\gamma_{n}f_{n}\in\Sigma[x].$ (8)
-Assuming the empty-set condition in 5 holds, a solution to 7 involving $\gamma_{1}(x)=1$ is guaranteed by Putinar’s P-satz to exist when the degrees of the polynomials are possibly unbounded.
+Assuming the empty-set condition in 6 holds, a solution to 8 involving $\gamma_{1}(x)=1$ is guaranteed by Putinar’s P-satz to exist when the degrees of the polynomials are possibly unbounded.
 In this case, the introduction of the SOS polynomial $\gamma_{1}(x)$ needlessly increases the set of feasible polynomials $f_{1}(x)$.
 However, when the degrees of the polynomials are truncated, this assertion may no longer hold true, justifying the utilization of $\gamma_{1}(x)$.
-For the rest of the paper, we refer to both inclusions, 6 and 7, as an SOS constraint.
+For the rest of the paper, we refer to both inclusions, 7 and 8, as an SOS constraint.
 III Specifications and Problem statement
 III-A Specifications
-The concept of a safety filter (see Figure 2) is based on the idea of enhancing the behavior of a previously designed legacy controller $u_{n}(x)$ with safety guarantees.
-The safety filter adjusts the control action of the legacy controller, ideally in a minimal way, to render the system safe with respect to an allowable set $\mathcal{X}_{a}$, encoding the state constraints, and a safe set $\mathcal{X}_{s}$.
-The safe set denotes a set of states that is forward invariant for some state feedback controller.
-Therefore, we require that the safe set is strictly contained within the allowable set, i.e.,
+The main objective of a safety filter (see Figure 2) is to enhance the behavior of a previously designed legacy controller $u_{n}(x)$ with safety guarantees.
+The safety filter adjusts the control action of the legacy controller, ideally in a minimal way, to ensure the system remains within a safe set $\mathcal{X}_{s}$, which is contained within an allowable set $\mathcal{X}_{a}$ that encodes the state constraints:
 $\mathcal{X}_{s}\subseteq\mathcal{X}_{a}.$ (9)
@@
 This nominal region $\mathcal{X}_{n}$ is required to be forward invariant under the legacy control $u_{n}(x)$, and to be strictly contained within the safe set, i.e.,
-$\mathcal{X}_{n}\subsetneq\mathcal{X}_{s}.$ (10)
-For a legacy controller that stabilizes the closed-loop system to an attractor within $\mathcal{X}_{a}$, a forward invariant set does always exist.
-Moreover, the advanced safety filter should preserve such a stability guarantee across the entire safe set.
-This is already accomplished within the nominal region $\mathcal{X}_{n}$ due to its forward invariance property.
-Hence, we only need to additionally ensure a finite-time convergence to $\mathcal{X}_{n}$ for all trajectories originating within the transitional region, defined as:
+$\mathcal{X}_{n}\subseteq\mathcal{X}_{s}.$ (11)
+For a legacy controller that stabilizes the closed-loop system to an attractor within $\mathcal{X}_{a}$, such a forward invariant set does always exist.
+Moreover, the advanced safety filter is designed to preserve the attractor of the legacy controller – assuming it lies within $\mathcal{X}_{a}$.
+This property is automatically satisfied within $\mathcal{X}_{n}$ due to the preservation property 10 and forward invariance of $\mathcal{X}_{n}$.
+Thus, it remains to ensure finite-time convergence of all trajectories originating outside $\mathcal{X}_{n}$ but within the safe set $\mathcal{X}_{s}$.
+We define the corresponding transitional region as:
 $\mathcal{X}_{t}:=\mathcal{X}_{s}\setminus\text{int}(\mathcal{X}_{n}),$ (12)
@@
 III-B CBF and CLF conditions formalizing the specifications
-Figure 3: The advanced safety filter is characterized by the following specifications: (a) the safe set $\mathcal{X}_{s}$ is forward invariant as encoded in 15, (b) it ensures finite-time convergence towards the nominal region $\mathcal{X}_{n}$ as encoded in 17, and (c) the nominal region is forward invariant under the legacy controller $u_{n}(x)$ as encoded in 19 assuming $d(x)=0$.
+Figure 3: The advanced safety filter is characterized by the following specifications: (a) the safe set $\mathcal{X}_{s}$ is forward invariant as encoded in 16, (b) it ensures finite-time convergence towards the nominal region $\mathcal{X}_{n}$ as encoded in 18, and (c) the nominal region is forward invariant under the legacy controller $u_{n}(x)$ as encoded in 20 assuming $d(x)=0$.
 We employ control Lyapunov and barrier functions to encode the specification from the previous subsection by introducing energy-like scalar functions.
-In the process, we define the safe set as the intersection of zero-sublevel sets of multiple continuously differentiable functions to closely fit the algebraic structure of the allowable set
-$\mathcal{X}_{a}:=\Bigl{\{}x\mid w_{i}(x)\leq 0\text{ }\forall i\in\mathcal{I}_{t}\Bigr{\}}=\bigcap_{i\in\mathcal{I}_{t}}\mathcal{X}_{a,i},$ (12)
-where $w_{i}\in R[x]$, $i\in\mathcal{I}_{t}$, are scalar polynomials, and their zero-sublevel sets $\mathcal{X}_{a,i}:=\{x\mid w_{i}(x)\leq 0\}$ are assumed to be compact.
-This compactness condition is important for applying Putinar’s P-satz in the following section of the paper, and it is not restrictive for the applications we have in mind.
-The forward invariance property of the safe set can be asserted by a scalar function $B(x)$ defined as the point-wise maximum of a set of continuously differentiable functions $B_{i}:\mathbb{R}^{n}\to\mathbb{R}$ indexed by $\mathcal{I}_{t}$:
-$B(x):=\underset{i\in\mathcal{I}_{t}}{\text{max}}\,B_{i}(x),$ (13)
+In the process, we define the safe set as the intersection of zero-sublevel sets of multiple continuously differentiable functions
+$\mathcal{X}_{a}:=\Bigl\{x\mid w_{i}(x)\leq 0\quad\forall i\in\mathcal{I}_{n_{\!B}}\Bigr\}=\bigcap_{i\in\mathcal{I}_{n_{\!B}}}\mathcal{X}_{a,i},$ (13)
+where $w_{i}\in R[x]$ are scalar polynomials, and $\mathcal{X}_{a,i}:=\{x\mid w_{i}(x)\leq 0\}$ denotes the corresponding zero-sublevel set.
+We assume that each set $\mathcal{X}_{a,i}$ is compact, which ensures that their intersection $\mathcal{X}_{a}$ is also compact.
+This compactness condition is important for the derivation of the SOS constraints, and it is well aligned with our intended application scenarios, where state constraints are typically bounded.
+The forward invariance property of the safe set can be asserted by a scalar function $B(x)$ defined as the point-wise maximum of a set of continuously differentiable functions $B_{i}:\mathbb{R}^{n}\to\mathbb{R}$ indexed by $\mathcal{I}_{n_{\!B}}$:
+$B(x):=\underset{i\in\mathcal{I}_{n_{\!B}}}{\text{max}}\,B_{i}(x),$ (14)
 such that the safe set is given by the intersection of the individual zero-sublevel sets:
-$\mathcal{X}_{s}:=\left\{x\mid B(x)\leq 0\right\}=\bigcap_{i\in\mathcal{I}_{t}}\mathcal{X}_{s,i},$
+$\mathcal{X}_{s}:=\left\{x\mid B(x)\leq 0\right\}=\bigcap_{i\in\mathcal{I}_{n_{\!B}}}\mathcal{X}_{s,i},$
 where $\mathcal{X}_{s,i}:=\left\{x\mid B_{i}(x)\leq 0\right\}$.
-Taking the intersection of multiple functions $B_{i}(x)$ allows for greater flexibility to increase the volume of the safe set compared to using the zero-sublevel set of a single CBF function.
-This is because the zero-sublevel sets of each $B_{i}(x)$ can extend beyond the allowable set, provided their intersection remains within the allowable set.
-In contrast, the zero-sublevel set of a single CBF must be fully contained within the allowable set.
-Hence, we encode the containment condition in 8 as:
+This multi-function construction provides greater flexibility in shaping the safe set compared to relying on a single barrier function.
+In particular, each function $B_{i}(x)$ is allowed to define a sublevel set that may partially extend beyond the allowable set, as long as their intersection $\mathcal{X}_{s}$ remains within the allowable set $\mathcal{X}_{a}$.
+By contrast, a single barrier function must define a sublevel set that is entirely contained within the allowable set, which leads to more conservative safe sets.
+Hence, to enforce the containment condition in 9, we impose the subset relation
 $\mathcal{X}_{s,i}\subseteq\mathcal{X}_{a,i},$ (15)
-for every $i\in\mathcal{I}_{t}$.
-This can be seen from $\bigcap_{i\in\mathcal{I}_{t}}\mathcal{X}_{s,i}\subseteq\bigcap_{i\in\mathcal{I}_{t}}\mathcal{X}_{a,i}$ implying $\mathcal{X}_{s}\subseteq\mathcal{X}_{a}$.
-Here, we assume that the CBF $B(x)$ is defined using the same number of polynomials $B_{i}(x)$ as polynomials $w_{i}(x)$ that define the allowable set.
-To reduce the number of decision variables in $B(x)$, we can equate certain polynomials $B_{i}(x)$ in the resulting SOS problem.
-This reduction results in a safe set comprising fewer distinct subsets $\mathcal{X}_{s,i}$ compared to the subsets $\mathcal{X}_{a,i}$ that define the allowable set.
+for every $i\in\mathcal{I}_{n_{\!B}}$.
+This condition ensures that the overall safe set $\mathcal{X}_{s}=\bigcap_{i\in\mathcal{I}_{n_{\!B}}}\mathcal{X}_{s,i}$ is contained within the allowable set $\mathcal{X}_{a}=\bigcap_{i\in\mathcal{I}_{n_{\!B}}}\mathcal{X}_{a,i}$.
+We assume that the barrier function $B(x)$ is constructed from the same number of components $B_{i}(x)$ as the number of polynomials $w_{i}(x)$ that define the allowable set.
+This condition is not restrictive in practice, since one can define multiple $B_{i}(x)$ functions to share the same expression, effectively reducing the number of distinct subsets $\mathcal{X}_{s,i}$ compared to the subsets $\mathcal{X}_{a,i}$.
+Conversely, it is also possible to define additional distinct $B_{i}(x)$ functions even if some $w_{i}(x)$ are repeated.
 Assumption 2
-We assume the safe set $\mathcal{X}_{s}$ to be regular in the sense that $\forall x\in\mathcal{X}_{s},\exists z\in\mathbb{R}^{n},\forall i\in\mathcal{I}_{t}$ [18, Definition 4.9]:
-$B_{i}(x)+\nabla B_{i}(x)^{T}z<0$
+We assume the safe set $\mathcal{X}_{s}$ to be regular in the sense that $\forall x\in\mathcal{X}_{s},\exists z\in\mathbb{R}^{n},\forall i\in\mathcal{I}_{n_{\!B}}$ [5, Def. 4.9]:
+$B_{i}(x)+\nabla B_{i}(x)^{\top}z<0$
 The regularity assumption in Assumption 2 is a mild one, and it generically holds true, if $B_{i}(x)$ is chosen as a polynomial.
-It implies that the interior of the safe set is given by $\text{int}(\mathcal{X}_{s})=\left\{x\mid B_{i}(x)<0\quad i\in\mathcal{I}_{t}\right\}$.
-Examples of polynomials that violate the regularity assumption can be found in [18, Chapter 4.2].
+It implies that the interior of the safe set is given by $\text{int}(\mathcal{X}_{s})=\left\{x\mid B_{i}(x)<0\,\,\forall i\in\mathcal{I}_{n_{\!B}}\right\}$.
+Examples of polynomials that violate the regularity assumption can be found in [5, Chapter 4.2].
 To ensure forward invariance of $\mathcal{X}_{s}$, as shown in Figure 3(a), $B(x)$ needs to satisfy Nagumo’s theorem, that is, $\forall x\in\partial\mathcal{X}_{s},\exists u\in\mathbb{R}^{m}$ (cf.
-[18, Nagumo’s Theorem 4.7]):
+[5, Nagumo’s Theorem 4.7]):
 $f(x)+G(x)u\in\mathcal{T}_{\mathcal{X}_{s}}(x),$ (16)
 where, under Assumption 2, the tangent cone is defined by
-$\mathcal{T}_{\mathcal{X}_{s}}(x):=\{z\mid\nabla B_{i}(x)^{T}z\leq 0\quad i\in\text{Act}(x)\},$ (16)
-and $\text{Act}(x):=\left\{i\in\mathcal{I}_{t}\mid B_{i}(x)=0\right\}$ denotes the set of active constraints.
-The function $B(x)$ in 13 satisfying 15 and 16 will henceforth be referred to as a Control Barrier Function (CBF).
-Compared to the standard safety filter, our approach also require finite-time convergence towards $\mathcal{X}_{n}$, as outlined in the specifications.
-This is asserted by a Control Lyapunov Function (CLF) defined by a continuously differentiable function $V:\mathbb{R}^{n}\to\mathbb{R}$.
-This function needs to satisfy a Control Lyapunov-like condition in the transitional region $\mathcal{X}_{t}$, as shown in Figure 3(b).
+$\mathcal{T}_{\mathcal{X}_{s}}(x):=\{z\mid\nabla B_{i}(x)^{\top}z\leq 0\quad i\in\text{Act}(x)\},$ (17)
+and $\text{Act}(x):=\left\{i\in\mathcal{I}_{n_{\!B}}\mid B_{i}(x)=0\right\}$ denotes the set of active constraints.
+The function $B(x)$ in 14 satisfying 16 and 17 will be referred to as a Control Barrier Function (CBF).
+Compared to the standard safety filter, our approach additionally require finite-time convergence to the nominal region $\mathcal{X}_{n}$.
+This is enforced via a Control Lyapunov Function (CLF), defined by a continuously differentiable function $V:\mathbb{R}^{n}\to\mathbb{R}$.
+In the transitional region $\mathcal{X}_{t}$, this function needs to satisfy a Control Lyapunov-like condition, as illustrated in Figure 3(b).
 Namely, $\forall x\in\mathcal{X}_{t},\exists u\in\mathbb{R}^{m}$:
-$\nabla V(x)^{T}\left(f(x)+G(x)u\right)+d(x)\leq 0$ (17)
+$\nabla V(x)^{\top}\left(f(x)+G(x)u\right)+d(x)\leq 0$ (18)
 with dissipation rate
@@
 where $\lambda:\mathbb{R}^{n}\to\mathbb{R}_{>0}$ is a Lipschitz continuous function, and $V_{0}\leq\text{min}_{x}\left(V(x)\right)$ is a constant value, ensuring $d(x)$ is strictly positive (and thus $V(x)$ strictly decaying) in the transitional region $\mathcal{X}_{t}$.
-By utilizing a strictly positive function $d(x)$, the CLF condition can be expressed as a non-strict inequality condition similar to the CBF condition – a crucial distinction for the subsequent arguments presented in the next section.
+By utilizing a strictly positive function $d(x)$, the CLF condition can be expressed as a non-strict inequality condition similar to the CBF condition – an important distinction for deriving the SOS constraints.
 We define the nominal region as the zero sublevel set of $V(x)$:
@@
 $\mathcal{X}_{t}:=\left\{x\mid B(x)\leq 0\leq V(x)\right\}.$
-The next lemma establishes the finite-time convergence property associated with the CLF $V(x)$. Its proof can be found in the appendix.
+The following lemma establishes the finite-time convergence property induced by the CLF $V(x)$. A detailed proof is provided in Section VIII-A .
 Lemma 1
-Given a Lipschitz-continuous controller $u_{s}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ that satisfies condition 17 with $u:=u_{s}(x)$ and renders the safe set $\mathcal{X}_{s}$ forward invariant, then for all trajectories of system 3 starting in the safe set $\mathcal{X}_{s}$ there exists some finite $T\geq 0$ such that $x(t)\in\mathcal{X}_{n}$ for all $t\geq T$.
+Given a Lipschitz-continuous controller $u_{s}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ that satisfies condition 18 with $u:=u_{s}(x)$ and renders the safe set $\mathcal{X}_{s}$ forward invariant, then for all trajectories of system $\dot{x}=f(x)+G(x)u_{s}(x)$ starting in the safe set $\mathcal{X}_{s}$ there exists some finite $T\geq 0$ such that $x(t)\in\mathcal{X}_{n}$ for all $t\geq T$.
 Additionally, to make the nominal region forward invariant w.r.t. the legacy controller $u_{n}(x)$, as depicted in Figure 3(c), we propose the condition that $\forall x\in\partial\mathcal{X}_{n}$:
-$\nabla V(x)^{T}\left(f(x)+G(x)u_{n}(x)\right)+d(x)\leq 0.$ (19)
+$\nabla V(x)^{\top}\left(f(x)+G(x)u_{n}(x)\right)+d(x)\leq 0.$ (20)
 Notice that forward invariance also follows with $d(x)=0$.
-By adopting a slightly stricter forward invariance condition in 19 involving the strictly positive dissipation rate $d(x)$, we ensure the existence of the safety filter implementation as specified in Section V.
+By adopting a slightly stricter forward invariance condition in 20 involving the strictly positive dissipation rate $d(x)$, we are able to proof the existence of a safety filter implementation in Section V.
 Remark 2
-As the volume of the transitional region $\mathcal{X}_{t}$ approaches zero, the proposed safety filter transitions into a safety filter based on a single CBF $V(x)=B_{1}(x)=\ldots=B_{t}(x)$, where the CBF and CLF condition 15 and 17 are automatically satisfied by the forward invariance condition 19 of the legacy controller.
-Therefore, feasible CBF and CLF candidates can be found (in the limit) if a forward invariant set for the system, governed by the legacy controller, exists and is fully contained within the allowable set.
+Compared to approaches that simultaneously establish safety and stability [17], our proposed specifications are less restrictive.
+In fact, feasible CBF and CLF candidates that satisfy these specifications can always be found, provided that a forward invariant set under the legacy controller exists – including a dissipation term as in 20 – that is fully contained within the allowable set.
+To such cases, a trivial but valid construction would be to set all barrier and Lyapunov functions equal, such that CBF and CLF condition 16 and 18 are automatically satisfied by the forward invariance condition of the legacy controller.
 III-C Problem statement
-The challenge of deriving the advanced safety filter can be divided into two subproblems.
-The first subproblem resolves around finding $B(x)$ and $V(x)$ that fulfill the CBF and CLF condition outlined in the preceding section.
-This involves constructing a rational controller $u_{\text{SOS}}(x)$ to prove the compatibility of the CLF and CBF conditions in 15 and 17.
-Furthermore, we necessitate the assurance of the forward invariance condition of the legacy controller as defined in 19.
+The advanced safety filter can be divided into two subproblems. The first resolves around finding $B(x)$ and $V(x)$ that fulfill the CBF and CLF condition outlined in the preceding section.
 Problem 1
 Given a dynamical system 3, an allowable set of states $\mathcal{X}_{a}$ and a legacy controller $u_{n}(x)\in\mathcal{U}$, find scalar functions $B(x):\mathbb{R}^{n}\to\mathbb{R}$ and $V(x):\mathbb{R}^{n}\to\mathbb{R}$ such that:
-  • (set containment) the containment conditions in 8 and 10 are satisfied,
-  • (finite-time convergence) the CLF condition in 17 is satisfied,
-  • (forward-invariance of $\mathcal{X}_{s}$) CBF and CLF conditions in 15 and 17 are satisfied on $\partial\mathcal{X}_{s}$ for the same $u$,
-  • (forward-invariance of $\mathcal{X}_{n}$) legacy control condition in 19 is satisfied,
+  • (set containment) the containment conditions in 9 and 11 are satisfied,
+  • (finite-time convergence) the CLF condition in 18 is satisfied,
+  • (forward-invariance of $\mathcal{X}_{s}$) CBF and CLF conditions in 16 and 18 are satisfied on $\partial\mathcal{X}_{s}$ for the same $u$,
+  • (forward-invariance of $\mathcal{X}_{n}$) legacy control condition in 20 is satisfied,
   • (objective) CBF $B(x)$ maximizes the volume of the safe set, denoted by $\text{vol}(\mathcal{X}_{s})$.
-For the second subproblem, a safety filter $u_{s}(x)$ is constructed from $B(x)$ and $V(x)$, which is guaranteed to be feasible and Lipschitz-continuous, entailing a smooth transitioning from the safety filter operation to the legacy control operation.
+For the second subproblem, a safety filter $u_{s}(x)$ is constructed based on the CBF and CLF.
 Problem 2
-Given a dynamical system 3, a Lipschitz-continuous legacy controller $u_{n}(x)$, a CBF $B(x)$ and CLF $V(x)$, as specified in Problem 1, find state-feedback control law $u_{s}(x)$ that is feasible for all $x\in\mathbb{R}^{n}$, satisfies the CBF and CLF conditions 15 and 17 with $u:=u_{s}(x)$, coincides with $u_{n}(x)$ on $\mathcal{X}_{n}$ (cf.
-9), and is Lipschitz-continuous.
+Given a dynamical system 3, a Lipschitz-continuous legacy controller $u_{n}(x)$, a CBF $B(x)$ and CLF $V(x)$, as specified in Problem 1, find state-feedback control law $u_{s}(x)$ that:
+  • is feasible for all $x\in\mathbb{R}^{n}$,
+  • satisfies the CBF and CLF conditions 16 and 18 with $u:=u_{s}(x)$,
+  • coincides with $u_{n}(x)$ on $\mathcal{X}_{n}$ (cf. 10),
+  • is Lipschitz-continuous.
 In Section IV, we delve into a solution for Problem 1, while a solution for Problem 2 is derived in Section V.
 Figure 4: Problem 1 involves searching for compatible CBF $B(x)$ and CLF $V(x)$ using SOS tools.
-Based on $B(x)$ and $V(x)$, the linear constraints of the QCQP given by $C(x)$, $b(x)$ and $r(x)$ are determined.
+Based on $B(x)$ and $V(x)$, the linear constraints of the QP given by $C(x)$, $b(x)$ and $r(x)$ are determined.
 IV Constructing CBF and CLF using SOS optimization
 IV-A SOS formulation of CBF and CLF conditions
-The objective of this subsection is to derive sufficient SOS conditions that encode the CBF condition on $\mathcal{X}_{s}$ and $\mathcal{X}_{n}$, as specified in 15 and 19, and the CLF condition, as specified in 17.
+The objective of this subsection is to derive sufficient SOS conditions that encode the CBF condition on $\mathcal{X}_{s}$ and $\mathcal{X}_{n}$, as specified in 16 and 20, and the CLF condition, as specified in 18.
 We further demonstrate that these SOS conditions are not overly conservative and pose no practical concerns.
-The CBF condition in 15 involves the set of active constraints $\text{Act}(x)$, a construction that cannot be directly expressed as an SOS constraint.
-To address this issue, we first consider the polynomials $B_{i}(x)$, $i\in\mathcal{I}_{t}$, defining the CBF $B(x)$ in 13, as independent CBFs in their own right.
-By ensuring compatibility among these CBFs $B_{i}(x)$ as proposed in [8], we demonstrate sufficiency of the resulting SOS constraints with respect to CBF condition in 15 (involving the desired CBF $B(x)$).
+The CBF condition in 16 involves the set of active constraints $\text{Act}(x)$, a construction that cannot be directly expressed as an SOS constraint.
+To address this issue, we first consider the polynomials $B_{i}(x)$, $i\in\mathcal{I}_{t}$, defining the CBF $B(x)$ in 14, as independent CBFs in their own right.
+By ensuring compatibility among these CBFs $B_{i}(x)$ as proposed in [18], we demonstrate sufficiency of the resulting SOS constraints with respect to CBF condition in 16 (involving the desired CBF $B(x)$).
 The CBF condition for each $B_{i}(x)$, $i\in\mathcal{I}_{t}$, is expressed as $\forall x\in\{x\in\mathcal{X}_{s}\mid B_{i}(x)=0\},\exists u\in\mathbb{R}^{m}$:
-$\nabla B_{i}(x)^{T}\left(f(x)+G(x)u\right)\leq 0.$ (20)
-Another challenge in converting conditions 20 and 17 to SOS constraints is the presence of the existential quantifier $\exists u\in\mathbb{R}^{m}$.
+$\nabla B_{i}(x)^{\top}\left(f(x)+G(x)u\right)\leq 0.$ (21)
+Another challenge in converting conditions 21 and 18 to SOS constraints is the presence of the existential quantifier $\exists u\in\mathbb{R}^{m}$.
 However, following the approach in [19], we can reformulate these conditions without the need for an existential quantifier.
-The CBF condition 20 can be expressed as $\forall x\in\{x\in\mathcal{X}_{s}\mid B_{i}(x)=0,\nabla B_{i}(x)^{T}G(x)=0\}$:
-$\nabla B_{i}(x)^{T}f(x)\leq 0.$ (21)
-Similarly, the CLF condition in 17 can be formulated as $\forall x\in\{x\in\mathcal{X}_{t}\mid\nabla V(x)^{T}G(x)=0\}$:
-$\nabla V(x)^{T}f(x)+d(x)\leq 0.$ (22)
+The CBF condition 21 can be expressed as $\forall x\in\{x\in\mathcal{X}_{s}\mid B_{i}(x)=0,\nabla B_{i}(x)^{\top}G(x)=0\}$:
+$\nabla B_{i}(x)^{\top}f(x)\leq 0.$ (22)
+Similarly, the CLF condition in 18 can be formulated as $\forall x\in\{x\in\mathcal{X}_{t}\mid\nabla V(x)^{\top}G(x)=0\}$:
+$\nabla V(x)^{\top}f(x)+d(x)\leq 0.$ (23)
 In order to render the problem tractable, we restrict our attention to polynomial candidates $B_{i}\in R[x]$, $i\in\mathcal{I}_{t}$, and polynomial candidate $V\in R[x]$.
-We then apply Putinar’s P-satz, as stated in Theorem 1, on a slightly more stringent conditions compared to 19, 21 and 22, where the inequality is replaced by a strict inequality.
+We then apply Putinar’s P-satz, as stated in Theorem 1, on a slightly more stringent conditions compared to 20, 22 and 23, where the inequality is replaced by a strict inequality.
 These more stringent conditions are summarized as the following empty-set conditions, $i\in\mathcal{I}_{t}$:
-$\begin{split}\{x\mid&-\nabla B_{i}(x)^{T}f(x)\leq 0,\nabla B_{i}(x)^{T}G(x)=0,\\ &B_{i}(x)=0,-B_{\mathcal{I}_{t}}(x)\geq 0\}=\emptyset\end{split}$ (23a)
-$\begin{split}\{x\mid&-\nabla V(x)^{T}f(x)-d(x)\leq 0,\nabla V(x)^{T}G(x)=0,\\ &V(x)\geq 0,-B_{\mathcal{I}_{t}}(x)\geq 0\}=\emptyset\end{split}$ (23b)
-$\begin{split}\{x\mid&\nabla V(x)^{T}\left(f(x)+G(x)u_{n}(x)\right)+d(x)\leq 0,\\ &V(x)=0\}=\emptyset,\end{split}$ (23c)
-where $B_{\mathcal{I}_{t}}(x)=\begin{bmatrix}B_{1}(x)&\ldots&B_{t}(x)\end{bmatrix}^{T}$ is a vector of CBFs.
+$\begin{split}\{x\mid&-\nabla B_{i}(x)^{\top}f(x)\leq 0,\nabla B_{i}(x)^{\top}G(x)=0,\\ &B_{i}(x)=0,-B_{\mathcal{I}_{t}}(x)\geq 0\}=\emptyset\end{split}$ (24a)
+$\begin{split}\{x\mid&-\nabla V(x)^{\top}f(x)-d(x)\leq 0,\nabla V(x)^{\top}G(x)=0,\\ &V(x)\geq 0,-B_{\mathcal{I}_{t}}(x)\geq 0\}=\emptyset\end{split}$ (24b)
+$\begin{split}\{x\mid&\nabla V(x)^{\top}\left(f(x)+G(x)u_{n}(x)\right)+d(x)\leq 0,\\ &V(x)=0\}=\emptyset,\end{split}$ (24c)
+where $B_{\mathcal{I}_{t}}(x)=\begin{bmatrix}B_{1}(x)&\ldots&B_{t}(x)\end{bmatrix}^{\top}$ is a vector of CBFs.
 The resulting SOS constraints are summarized below.
-A detailed derivation of the SOS constraint 24 from the empty-set condition 23 is given in the supplementary material in [19].
+A detailed derivation of the SOS constraint 25 from the empty-set condition 24 is given in Section VIII-B.
 $\begin{split}-\nabla B_{i}^{T}\left(s_{i}f+Gp_{i}\right)&\in\textbf{M}(B_{i},-B_{\mathcal{I}_{t}})\end{split}$ (25a)
@@
 For clarity, we omit the brackets and variables in the function notation.
-To achieve the simultaneous satisfaction of CBF and CLF conditions 15 and 17 with a single input $u:=p(x)/s(x)$ on $\partial\mathcal{X}_{s}$, as specified in Problem 1, we equate the polynomials $p_{i}(x)$ and $s_{i}(x)$, $i\in\{0\}\cup\mathcal{I}_{t}$.
+To achieve the simultaneous satisfaction of CBF and CLF conditions 16 and 18 with a single input $u:=p(x)/s(x)$ on $\partial\mathcal{X}_{s}$, as specified in Problem 1, we equate the polynomials $p_{i}(x)$ and $s_{i}(x)$, $i\in\{0\}\cup\mathcal{I}_{t}$.
 The resulting SOS constraints are given as:
@@
 Remark 3
-The polynomials $p(x)$ and $s(x)$ induce (after dividing 25a and 25b by $s(x)$) a rational controller $u_{\text{SOS}}(x):=p(x)/s(x)$, employed to prove compatibility of the CBFs $B_{i}(x)$ and CLF $V(x)$.
-The proof of Lemma 2 can be found in the appendix.
+The polynomials $p(x)$ and $s(x)$ induce (after dividing 26a and 26b by $s(x)$) a rational controller $u_{\text{SOS}}(x):=p(x)/s(x)$, employed to prove compatibility of the CBFs $B_{i}(x)$ and CLF $V(x)$.
+The proof of Lemma 2 can be found in Section VIII-C.
 Lemma 2
-If $s(x)>0$ and Assumption 2 is satisfied, then the SOS constraints in 25 are sufficient conditions for the CBF and CLF property in 15, 17 and 19 to hold simultaneously with $u:=p(x)/s(x)$.
-The SOS constraints in 25 imply the inequalities since $p\in\Sigma[x]$ implies $p(x)\geq 0$:
-$\begin{split}\nabla B_{i}^{T}\left(sf+Gp\right)+\gamma_{0,i}B_{i}-\gamma_{1,i}^{T}B_{\mathcal{I}_{t}}&\leq 0\end{split}$ (26a)
-$\begin{split}\nabla V^{T}\left(sf+Gp\right)+sd+\gamma_{0,0}V-\gamma_{1,0}^{T}B_{\mathcal{I}_{t}}&\leq 0\end{split}$ (26b)
-$\begin{split}\nabla V^{T}\left(f+Gu_{n}\right)+d+\gamma_{n}V&\leq 0,\end{split}$ (26c)
-for some $\gamma_{0,i},\gamma_{n}\in R[x]$, $p\in(R[x])^{m}$, $s,\gamma_{0,0}\in\Sigma[x]$, and $\gamma_{1,i},\gamma_{1,0}\in(\Sigma[x])^{t}$, $i\in\mathcal{I}_{t}$.
+If $s(x)>0$ and Assumption 2 is satisfied, then the SOS constraints in 26 are sufficient conditions for the CBF and CLF property in 16, 18 and 20 to hold simultaneously with $u:=p(x)/s(x)$.
+The SOS constraints in 26 imply the inequalities since $p\in\Sigma[x]$ implies $p(x)\geq 0$:
+$\begin{split}\nabla B_{i}^{\top}\left(sf+Gp\right)+\gamma_{0,i}B_{i}-\gamma_{1,i}^{\top}B_{\mathcal{I}_{t}}&\leq 0\end{split}$ (27a)
+$\begin{split}\nabla V^{\top}\left(sf+Gp\right)+sd+\gamma_{0,0}V-\gamma_{1,0}^{\top}B_{\mathcal{I}_{t}}&\leq 0\end{split}$ (27b)
+$\begin{split}\nabla V^{\top}\left(f+Gu_{n}\right)+d+\gamma_{n}V&\leq 0,\end{split}$ (27c)
+for some $\gamma_{0,i},\gamma_{n}\in R[x]$, $p\in(R[x])^{m}$, $s,\gamma_{0,0}\in\Sigma[x]$, and $\gamma_{1,i},\gamma_{1,0}\in(\Sigma[x])^{\top}$, $i\in\mathcal{I}_{t}$.
 Lemma 2 only provides sufficient conditions.
-However, adding an arbitrary small perturbation $\epsilon>0$ to the SOS constraints in 25 results in necessary conditions when restricting to rational controllers $u_{\text{SOS}}(x)=p(x)/s(x)$.
-This is because 15, 17 and 19 with $u:=u_{\text{SOS}}(x)$ imply strict inequalities involving $\epsilon$ as follows: $p(x)\leq 0$ implies $p(x)-\epsilon<0$.
-These strict inequalities imply the SOS constraints 25 perturbed by $\epsilon$ according to Theorem 1.
-Hence, for all practical purposes, the solution set of the SOS constraints in 25 encompasses the desired set of CBFs and CLFs.
+However, adding an arbitrary small perturbation $\epsilon>0$ to the SOS constraints in 26 results in necessary conditions when restricting to rational controllers $u_{\text{SOS}}(x)=p(x)/s(x)$.
+This is because 16, 18 and 20 with $u:=u_{\text{SOS}}(x)$ imply strict inequalities involving $\epsilon$ as follows: $p(x)\leq 0$ implies $p(x)-\epsilon<0$.
+These strict inequalities imply the SOS constraints 26 perturbed by $\epsilon$ according to Theorem 1.
+Hence, for all practical purposes, the solution set of the SOS constraints in 26 encompasses the desired set of CBFs and CLFs.
 IV-B SOS formulation of input constraints
@@
 $\begin{split}-a_{i}p(x)+b_{i}s(x)\in\textbf{M}(-B_{\mathcal{I}_{t}}),\end{split}$ (29)
-where $a_{i}$ is the $i$th row of $A$, $i\in\mathcal{I}_{l}$, and $p(x)$ and $s(x)$ are as in 25. Next, we consider the quadratic input constraint
-$\mathcal{U}=\{u\mid(u-u_{0})^{T}Q_{u}(u-u_{0})\leq u_{\text{max}}^{2}\},$ (29)
-for some $Q_{u}\succeq 0$ and $u_{0}\in\mathbb{R}^{m}$. The quadratic input constraint 29 can be encoded using the SOS constraint involving variables $x\in\mathbb{R}^{n}$ and $v\in\mathbb{R}^{m}$:
-$\begin{split}&-v^{T}Q_{u}\left(p(x)-s(x)u_{0}\right)+s(x)u_{\text{max}}^{2}\in\textbf{M}(-f_{u},-B_{\mathcal{I}_{t}}),\end{split}$ (30)
-where $f_{u}(x,v):=v^{T}Q_{u}v-u_{\text{max}}^{2}\in R[x,v]$ encodes the input constraint set in 29.
+where $a_{i}$ is the $i$th row of $A$, $i\in\mathcal{I}_{l}$, and $p(x)$ and $s(x)$ are as in 26. Next, we consider the quadratic input constraint
+$\mathcal{U}=\{u\mid(u-u_{0})^{\top}Q_{u}(u-u_{0})\leq u_{\text{max}}^{2}\},$ (30)
+for some $Q_{u}\succeq 0$ and $u_{0}\in\mathbb{R}^{m}$. The quadratic input constraint 30 can be encoded using the SOS constraint involving variables $x\in\mathbb{R}^{n}$ and $v\in\mathbb{R}^{m}$:
+$\begin{split}&-v^{\top}Q_{u}\left(p(x)-s(x)u_{0}\right)+s(x)u_{\text{max}}^{2}\in\textbf{M}(-f_{u},-B_{\mathcal{I}_{t}}),\end{split}$ (31)
+where $f_{u}(x,v):=v^{\top}Q_{u}v-u_{\text{max}}^{2}\in R[x,v]$ encodes the input constraint set in 30.
 The following lemma establishes the implication of the input constraints based on the SOS constraints.
-The proof is reported in the appendix.
+The proof is given in Section VIII-D.
 Lemma 3
-If $s(x)>0$, a solution $p(x)$ and $s(x)$ to the SOS constraint 28, resp. 30, also satisfies the input constraints 27, resp. 29, with $p(x)/s(x)\in\mathcal{U}$ for all $x\in\mathcal{X}_{s}$.
+If $s(x)>0$, a solution $p(x)$ and $s(x)$ to the SOS constraint 29, resp. 31, also satisfies the input constraints 28, resp. 30, with $p(x)/s(x)\in\mathcal{U}$ for all $x\in\mathcal{X}_{s}$.
 IV-C Safe set and nominal region requirement
-In this section, we express the containment conditions of the allowable set $\mathcal{X}_{a}$, the safe set $\mathcal{X}_{s}$ and the nominal region $\mathcal{X}_{n}$, as defined in 14 and 10 (cf.
+In this section, we express the containment conditions of the allowable set $\mathcal{X}_{a}$, the safe set $\mathcal{X}_{s}$ and the nominal region $\mathcal{X}_{n}$, as defined in 15 and 11 (cf.
 Figure 1), in terms of SOS constraints.
-A slightly more stringent condition compared to 14, can be expressed as the empty-set condition for all $i\in\mathcal{I}_{t}$:
+A slightly more stringent condition compared to 15, can be expressed as the empty-set condition for all $i\in\mathcal{I}_{t}$:
 $\{x\mid B_{i}(x)\leq 0,w_{i}(x)\geq 0\}=\emptyset.$ (32)
-Using Putinar’s P-satz in Theorem 1, we translate 31 into the SOS constraints for all $i\in\mathcal{I}_{t}$:
+Using Putinar’s P-satz in Theorem 1, we translate 32 into the SOS constraints for all $i\in\mathcal{I}_{t}$:
 $B_{i}\in\textbf{M}(w_{i}).$ (33)
-Similarly, the containment condition of $\mathcal{X}_{n}$ inside $\mathcal{X}_{s}$ in 10, can be expressed as the empty-set condition for all $i\in\mathcal{I}_{t}$:
+Similarly, the containment condition of $\mathcal{X}_{n}$ inside $\mathcal{X}_{s}$ in 11, can be expressed as the empty-set condition for all $i\in\mathcal{I}_{t}$:
 $\{x\mid V(x)\leq 0,B_{i}(x)\geq 0\}=\emptyset.$ (34)
-Using Putinar’s P-satz, we translate 33 into the SOS constraints for all $i\in\mathcal{I}_{t}$:
+Using Putinar’s P-satz, we translate 34 into the SOS constraints for all $i\in\mathcal{I}_{t}$:
 $-B_{i}\in\textbf{M}(V).$ (35)
 IV-D Resulting SOS problem
-The solution to Problem 1, i.e., finding suitable functions $B_{i}(x)$, $i\in\mathcal{I}_{t}$, and $V(x)$ can be obtained by solving the following SOS optimization problem:
+The solution to Problem 1 – that is the synthesis of suitable functions $B_{i}(x)$, $i\in\mathcal{I}_{n_{\!B}}$, and $V(x)$ – can be obtained by solving the following SOS optimization problem:
 […]
-where 25 encode the CBF and CLF condition, 28 and 30 encode linear and quadratic input constraints, 32 and 34 encode the set containment properties, and a constant $\epsilon>0$ assures that $s(x)>0$.
+Here, the constraints encode set the CBF and CLF conditions, the containment properties 9 and 11, and a constant $\epsilon>0$ assures that $s(x)>0$.
+Additional SOS constraints can be incorporated to capture linear and quadratic input constraint 29 and 31.
 IV-E Alternating Algorithm
-The solution of the SOS problem 39 can be found by iteratively alternate between searching over one set of decision variables while keeping the others fixed, see [11, 8].
+The solution of the SOS problem 40 can be found by iteratively alternate between searching over one set of decision variables while keeping the others fixed, see [19, 18].
 In this process, either constraint margins are minimized or the volume of $\mathcal{X}_{s}$ is maximized.
-The volume is approximated by the summation of traces of $Q_{B_{i}}$ in $B_{i}(x)=Z_{B,i}(x)^{T}Q_{B,i}Z_{B,i}(x)$, similar to the approach in [7, 20].
+The volume is approximated by the summation of traces of the gram matrix $Q_{B_{i}}$ in $B_{i}(x)=Z_{B,i}(x)^{\top}Q_{B,i}Z_{B,i}(x)$, similar to the approach in [23, 12].
 Such an alternating algorithm requires a feasible initialization of the decision variables that are kept fixed in the first iteration.
 This is achieved through an initialization procedure that starts with an initially empty region of interest and expands it until it encompasses the allowable set $\mathcal{X}_{a}$.
-For more details, we refer the reader to Section IX.
+For more details, we refer the reader to the supplementary material of this paper.
 V Safety Filter Implementation
-The objective of this section is to implement the advanced safety filter as specified in Problem 2.
-Given a linear and quadratic input constraints set $\mathcal{U}$, we propose the following realization of the Quadratically Constrained Quadratic Program (QCQP) in closed loop for every state $x\in\mathbb{R}^{n}$ along the trajectory:
-$\begin{array}[]{llll}u_{s}(x):=&\underset{u\in\mathcal{U}}{\mbox{min}}&\|u_{n}(x)-u\|^{2}\\ &\mbox{s.t.}&C(x)u+b(x)\leq r(x),\\ \end{array}$ (42)
+Given a linear input constraints set $\mathcal{U}$, we propose the following realization of the QP in closed loop for every state $x\in\mathbb{R}^{n}$ along the trajectory:
+[…]
 where
-$C(x)$ $:=\begin{bmatrix}\nabla V(x)^{T}G(x)\\ \nabla B_{1}(x)^{T}G(x)\\ \vdots\\ \nabla B_{t}(x)^{T}G(x)\\ \end{bmatrix}$ (43a)
-$b(x)$ $:=\begin{bmatrix}\nabla V(x)^{T}f(x)+d(x)\\ \nabla B_{1}(x)^{T}f(x)\\ \vdots\\ \nabla B_{t}(x)^{T}f(x)\end{bmatrix}$ (43b)
-encode the CLF and CBF conditions 15 and 17.
-The smooth functions $r(x)=\begin{bmatrix}r_{1}(x)&\ldots&r_{t}(x)\end{bmatrix}$ are state-dependent slack variables.
-Within the domain of CLF and CBF conditions 15 and 17, they must satisfy the upper bounds, given by:
+$C(x):=\left[\begin{smallmatrix}\nabla V(x)^{\top}G(x)\\ \nabla B_{1}(x)^{\top}G(x)\\ \vdots\\ \nabla B_{n_{\!B}}(x)^{\top}G(x)\\ \end{smallmatrix}\right],\,b(x):=\left[\begin{smallmatrix}\nabla V(x)^{\top}f(x)+d(x)\\ \nabla B_{1}(x)^{\top}f(x)\\ \vdots\\ \nabla B_{n_{\!B}}(x)^{\top}f(x)\end{smallmatrix}\right]$ (44a)
+encode the CLF and CBF conditions 16 and 18.
+The smooth functions $r(x)=[\begin{matrix}r_{1}(x)&\ldots&r_{n_{\!B}}(x)\end{matrix}]$ represent state-dependent slack variables that specify the allowed dissipation rate and safety violation.
+These slack-variables can be found using an SOS optimization, subject to upper and lower bounds that enforce the specifications outlined in Problem 2.
+Remark 4
+In the literature on safety filters, see for example [1], the negative of the allowed dissipation rate and the safety violation are usually specified by class $\mathcal{K}$ functions $\alpha(V(x))$ and $\gamma_{i}(B_{i}(x))$:
+$\nabla V(x)^{\top}(f(x)+G(x)u)\leq-\alpha(V(x)),$ (45a)
+$\nabla B_{i}(x)^{\top}(f(x)+G(x)u)\leq-\gamma_{i}(B_{i}(x))\quad i\in\mathcal{I}_{n_{\!B}}.$ (45b)
+In this case, the class $\mathcal{K}$ functions $\alpha$ and $\gamma_{i}$ depend on the value of $V(x)$ and $B_{i}(x)$.
+However, in 44, we consider a broader class of functions $r_{0}(x)$ and $r_{i}(x)$, $i\in\mathcal{I}_{n_{\!B}}$, to ensure the preservation of the legacy control behavior during nominal conditions.
+V-A Upper bounds on $r(x)$
+Within the domain of CLF and CBF conditions 16 and 18, the slack-variables must satisfy the upper bounds, given by
 $r_{0}(x)\leq 0$ (46)
@@
 $r_{i}(x)\leq 0$ (47)
-for all $i\in\mathcal{I}_{t}$ and $x\in\partial\mathcal{X}_{s}$.
-On the one hand, selecting a larger positive value for $r_{0}(x)$ within $\mathcal{X}_{n}$ (i.e.
-outside the domain of the CLF condition) increases the allowed dissipation rate violation in terms of larger values of $\dot{V}(x)=\nabla V(x)^{T}\dot{x}$.
-On the other hand, selecting a larger positive value for $r_{i}(x)$, $i\in\mathcal{I}_{t}$, within $\mathcal{X}_{s}$ (i.e.
-outside the domain of the CBF condition) increases the allowed safety violation in terms of larger positive values of $\dot{B}(x)=\nabla B(x)^{T}\dot{x}$.
-This renders the safety filter more passive within the interior of the safe set, and thus leaves more room to satisfy the CLF constraint.
-However, an overly large value for $r_{i}(x)$, $i\in\mathcal{I}_{t}$, can result in abrupt modifications of the control action when approaching the boundary of the safe set.
-Remark 4
-In the literature on safety filters, see for example [4], the negative of the allowed dissipation rate and the safety violation are usually specified by class $\mathcal{K}$ functions $\alpha(V(x))$ and $\gamma_{i}(B_{i}(x))$:
-$\nabla V(x)^{T}(f(x)+G(x)u)\leq-\alpha(V(x)),$ (46a)
-$\nabla B_{i}(x)^{T}(f(x)+G(x)u)\leq-\gamma_{i}(B_{i}(x))\quad i\in\mathcal{I}_{t}.$ (46b)
-In this case, the class $\mathcal{K}$ functions $\alpha$ and $\gamma_{i}$ depend on the value of $V(x)$ and $B_{i}(x)$.
-However, in 43, we consider a broader class of the allowed dissipation rate violation $r_{0}(x)$ and allowed safety violation, $r_{i}(x)$, $i\in\mathcal{I}_{t}$, to ensure a smooth safety filter behavior, while satisfying further requirements presented below.
-To ensure perfect tracking of the legacy controller $u_{n}(x)$ within $\mathcal{X}_{n}$ (cf. 9), we further lower bound $r(x)$ as:
+for all $i\in\mathcal{I}_{n_{\!B}}$ and $x\in\partial\mathcal{X}_{s}$.
+V-B Lower bounds on $r(x)$
+To ensure tracking of the legacy controller $u_{n}(x)$ within $\mathcal{X}_{n}$ (cf. 10), we impose the following lower bound on $r(x)$:
 $C(x)u_{n}(x)+b(x)\leq r(x)$ (48)
 for all $x\in\mathcal{X}_{n}$.
-Indeed, condition 47 ensures that the cost of the QCQP in 42 can be minimized to zero within $\mathcal{X}_{n}$ with $u=u_{n}(x)$, making $u_{s}(x)$ coincide with $u_{n}(x)\in\mathcal{U}$.
-To ensure existence of $r(x)$ satisfying lower and upper bounds 44, 45 and 47, we need to show for $r_{0}(x)$ that $\nabla V(x)^{T}\left(f(x)+G(x)u_{n}(x)\right)+d(x)\leq 0$ on the intersection of $\mathcal{X}_{t}$ (cf.
-44) and $\mathcal{X}_{n}$ (cf.
-47).
-This is however directly ensured by forward invariance condition 19.
-Furthermore, we need to show for $r_{i}(x)$, $i\in\mathcal{I}_{t}$, that $\nabla B_{i}(x)^{T}\left(f(x)+G(x)u_{n}(x)\right)\leq 0$ on the intersection of $\partial\mathcal{X}_{s}$ (cf.
-45) and $\mathcal{X}_{n}$ (cf.
-47).
-This is however automatically satisfied as the domains do not overlap due to the containment condition 10.
-Additionally, the selection of $r(x)$ must be lower-bounded to ensure the feasibility of the QCQP 42 within $\mathcal{X}_{t}$, as 47 already ensures feasibility within $\mathcal{X}_{n}$.
-Although the exact lower bound is unknown, it can be approximated using a feasible (but not necessarily optimal) solution $u_{\text{sos}}(x)$ from Section IV:
-$C(x)u_{\text{sos}}(x)+b(x)\leq r(x),$ (48)
-for all $x\in\mathcal{X}_{t}$. To guarantee the existence of such an $r(x)$, note that the lower bound 48 satisfies the upper bound conditions 44 and 45 as a result of the SOS constraints 25a and 25b.
+This condition guarantees that the cost in the QP in 43 can be minimized to zero within $\mathcal{X}_{n}$ by choosing $u=u_{n}(x)$, ensuring that the safety filter $u_{s}(x)$ coincides with the legacy control action $u_{n}(x)\in\mathcal{U}$.
+The existence of $r_{0}(x)$ satisfying 46 and 48 is ensured through 20, while the existence of $r_{i}(x)$, $i\in\mathcal{I}_{n_{\!B}}$, satisfying 47 and 48 is ensured if $\mathcal{X}_{n}\subseteq\text{int}(\mathcal{X}_{s}$).
+Furthermore, the selection of $r(x)$ must be lower-bounded to ensure the feasibility of the QP 43 within $\mathcal{X}_{s}$.
+Although the exact lower bound is unknown, it can not be directly obtain as it involves an existence quantifier $\exists u\in\mathcal{U}$ for every state $x$.
+However, it can be approximated using a feasible (but not necessarily optimal) solution $u_{\text{sos}}(x)$ from Section IV:
+$C(x)u_{\text{sos}}(x)+b(x)\leq r(x)$ (49)
+for all $x\in\mathcal{X}_{t}$.
+The existence of $r(x)$ satisfying 46, 47 and 49 is guaranteed through the SOS constraints 26a and 26b.
+Typical lower and upper bound selections for $r_{0}(x)$ are illustrated in the supplementary material of this paper.
+V-C Resulting SOS Problem
+The polynomial candidate $r(x)$ satisfying 46, 47, 48 and 49 can be obtained by solving the SOS problem:
+[…]
+where $B_{i}$, $V$, $p$, $s$ are taken from solving the SOS problem in 40, and the gram matrix $Q_{r,i}$ is defined by $r_{i}(x)=Z(x)^{\top}Q_{r,i}Z(x)$.
+The cost function in 58 favors choices of $r(x)$ that yield smooth intervention of the safety filter as the system approaches the state constraints.
+V-D Lipschitz-continuity
 Finally, to establish local Lipschitz-continuity of the proposed safety filter, we need the following assumption.
 Assumption 3 (LICQ)
-The Linear Independence Constraint Qualification (LICQ) is satisfied for the QCQP in 43 across all states $x$ and inputs $u$, that is, the gradients w.r.t.
-$u$ of the active inequality constraints in 43 are linearly independent.
+The Linear Independence Constraint Qualification (LICQ) is satisfied for the QP in 44 across all states $x$ and inputs $u$, that is, the gradients w.r.t.
+$u$ of the active inequality constraints in 44 are linearly independent.
 Equivalently, the rows of $C(x)$ associated with the active inequality constraints are linearly independent.
 Remark 5
-By extending the QCQP in 43 with a slack variable $\delta\in\mathbb{R}^{t+1}$ as in [21]
-$\begin{array}[]{llll}u_{s}(x):=&\underset{(u,\delta)\in\mathcal{U}\times\mathbb{R}^{t+1}}{\mbox{argmin}}&\|u_{n}(x)-u\|^{2}+\lambda\|\delta\|^{2}\\ &\mbox{s.t.}&C(x)^{T}u+b(x)\leq\delta,\\ \end{array}$
+By extending the QP in 44 with a slack variable $\delta\in\mathbb{R}^{t+1}$ as in [3]
+[…]
 we automatically satisfy the LICQ property, as demonstrated in [3].
@@
 [10][Theorem 3.1]), but can also avoid the resulting feedback controller from chattering, as explained in [15].
-The following theorem establishes local Lipschitz-continuity of the developed QCQP controller.
-Its proof can be found in the appendix.
+The following theorem establishes local Lipschitz-continuity of the developed QP controller.
+Its proof is given in Section VIII-E.
 Theorem 2
-Under Assumption 3, given a smooth $B(x)$ and $V(x)$ as specified in Problem 1, a smooth legacy controller $u_{n}(x)$, and twice differentiable $r_{i}(x)$, $i\in\{0\}\cup\mathcal{I}_{t}$, lower bounded by 48, then $u_{s}(x)$ as defined in 43 is locally Lipschitz-continuous on the safe set $\mathcal{X}_{s}$.
-We can find polynomial candidates $r(x)$ based on given CBF $B(x)$ and CLF $V(x)$ by solving an additional SOS problem that encodes the upper and lower bounds 44, 45, 47 and 48.
+Under Assumption 3, given a smooth $B(x)$ and $V(x)$ as specified in Problem 1, a smooth legacy controller $u_{n}(x)$, and twice differentiable $r_{i}(x)$, $i\in\{0\}\cup\mathcal{I}_{n_{\!B}}$, lower bounded by 49, then $u_{s}(x)$ as defined in 44 is locally Lipschitz-continuous on the safe set $\mathcal{X}_{s}$.
 VI Power Converter Case Study
@@
 Figure 5: The schematic is presented in the (dq) rotating frame, compromising a DC-link, 3-phase bridge, and an ac-line connected to an infinite bus $v_{g}$.
-Assuming the average-value model for the 3-phase bridge, the input $\bar{u}=\begin{bmatrix}m_{d}&m_{q}\end{bmatrix}^{T}$ determines the converter voltage $v_{c}$ as a function of the dc-link voltage $v_{dc}$.
-Likewise, the dc-current flowing into the capacitor $c_{dc}$ is given by the input $\bar{u}$ and the ac-line current $i=\begin{bmatrix}i_{d}&i_{q}\end{bmatrix}^{T}$.
+Assuming the average-value model for the 3-phase bridge, the input $\bar{u}=\begin{bmatrix}m_{d}&m_{q}\end{bmatrix}^{\top}$ determines the converter voltage $v_{c}$ as a function of the dc-link voltage $v_{dc}$.
+Likewise, the dc-current flowing into the capacitor $c_{dc}$ is given by the input $\bar{u}$ and the ac-line current $i=\begin{bmatrix}i_{d}&i_{q}\end{bmatrix}^{\top}$.
 The states consisting of the dc-link voltage $v_{dc}$ and the line-current $i$ are measured.
@@
 consisting of transformer and line inductance, $c_{dc}=0.02$ is the dc-link capacitance in p.u., and $g_{dc}:=1/r_{dc}=0.001$ is the admittance of the DC-link capacitor in p.u.
-The measured states are denoted by $\bar{x}:=\begin{bmatrix}v_{dc}&i_{d}&i_{q}\end{bmatrix}^{T}$, and the input is given by $\bar{u}:=\begin{bmatrix}m_{d}&m_{q}\end{bmatrix}^{T}$, describing the modulation indices.
-The objective is to stabilize both the DC-link voltage $v_{dc}$ to $1$, and the quadrature current $i_{q}$ to its reference $i_{q,ref}$, which is assumed to be constant.
-As a consequence, to meet the steady-state specifications, the direct current $i_{d}$ is stabilized to $-g_{dc}$.
+The measured states are denoted by $\bar{x}:=[\begin{matrix}v_{dc}&i_{d}&i_{q}\end{matrix}]^{\top}$, and the input is given by $\bar{u}:=[\begin{matrix}m_{d}&m_{q}\end{matrix}]^{\top}$, describing the modulation indices.
+The objective is to stabilize both the DC-link voltage $v_{dc}$ to 11, and the quadrature current $i_{q}$ to its reference $i_{q,ref}$, which is assumed to be constant.
+As a consequence, to meet the steady-state specifications, the direct current $i_{d}$ is stabilized to $g_{dc}$.
 The allowable set of states must adhere to the following inequality constraints:
 $0.2\leq v_{dc}\leq 1.2,\quad i_{d}^{2}+i_{q}^{2}\leq(1.2)^{2}.$ (60)
-Furthermore, the quadratic input constraint $\mathcal{U}$ is given by
-$m_{d}^{2}+m_{q}^{2}\leq(1.2)^{2}.$ (51)
 The system’s equilibrium is characterized by
 $x^{*}=\begin{bmatrix}1\\ -g_{dc}\\ i_{q,ref}\end{bmatrix},u^{*}=\begin{bmatrix}1-\omega li_{q,ref}\\ -\omega lg_{dc}\end{bmatrix}$.
-We introduce the error coordinates for the states $x=\bar{x}-x^{*}=\begin{bmatrix}\tilde{v}_{dc}&\tilde{i}_{d}&\tilde{i}_{q}\end{bmatrix}^{T}$ and the input variables $u=\bar{u}-u^{*}=\begin{bmatrix}\tilde{m}_{d}&\tilde{m}_{q}\end{bmatrix}^{T}$ such that $\dot{x}=0$ at $x=0$ and $u=0$.
-The system represented in error coordinates is described by system 2 with
+We introduce the error coordinates for the states $x=\bar{x}-x^{*}=[\begin{matrix}\tilde{v}_{dc}&\tilde{i}_{d}&\tilde{i}_{q}\end{matrix}]^{\top}$ and the input variables $u=\bar{u}-u^{*}=[\begin{matrix}\tilde{m}_{d}&\tilde{m}_{q}\end{matrix}]^{\top}$ such that $\dot{x}=0$ at $x=0$ and $u=0$.
+The system represented in error coordinates is described by system $\dot{x}=f(x)+G(x)u$ with
 $f(x)=\begin{bmatrix}-0.05\tilde{v}_{dc}-57.9\tilde{i}_{d}+0.00919\tilde{i}_{q}\\ 1710\tilde{v}_{dc}+314\tilde{i}_{q}\\ -0.271\tilde{v}_{dc}-314\tilde{i}_{d}\end{bmatrix}$
@@
 In the following subsection, we first consider a single quadrature current reference set to $i_{q,ref}=0$.
-Then, we introduce a quadrature current reference as a new state to the dynamical system, with a zero time-derivative, i.e., ${\dot{i}}_{q,ref}=0$.
-For each scenario, we find the CBFs $B_{i}(x)$ and CLF $V(x)$ (cf.
-Problem 1), and perform simulations that illustrate the additional guarantees provided by the advanced safety filter $u_{s}(x)$ compared with a standard safety filter implementation.
-We use CVXOPT Python library to solve the SDPs of the alternating algorithm, and to implement the QCQP in 42.
-This library implements an optimization problem solver based on second-order and positive semidefinite cones.
-The simulations are performed using the ODE solver from scipy with a sampling time of $1$ $\mu s$.
-The full code used for simulations and optimization is publicly available online [^1] https://github.com/MichaelSchneeberger/advanced-safety-filter/.
+Then, we introduce a non-zero quadrature current reference as a new state to the dynamical system.
+The full code used for the optimization is available online [^1] https://github.com/MichaelSchneeberger/advanced-safety-filter/.
 Figure 6: The allowable set $\mathcal{X}_{a}$, the safe set $\mathcal{X}_{s}$, and the nominal region $\mathcal{X}_{n}$ projected to the $\tilde{v}_{dc}$ and $\tilde{i}_{d}$ coordinate are shown for $\tilde{i}_{q}=i_{q,ref}=0$.
-In addition, the zero-level sets of the polynomials $B_{1}(x)$ and $B_{2}(x)$ are shown in grey.
+In addition, the zero-level sets of the polynomials $B_{1}(x)$ and $B_{2}(x)$ are shown in gray.
 Three scenarios are shown, where (a) the input action is subjected to no constraints, i.e.
-$u(x(t))\in\mathbb{R}^{m}$, (b) the input action is subjected to a quadratic constraint given by $u(x(t))\in\mathcal{U}:=\{u\mid u^{T}u\leq 1.3\}$, and (c) $i_{q,ref}$ is integrated as a new state of the dynamical system using the same input constraints.
+$u(x(t))\in\mathbb{R}^{m}$, (b) the input action is subjected to a quadratic constraint given by $u(x(t))\in\mathcal{U}:=\{u\mid u^{\top}u\leq 1.2\}$, and (c) $i_{q,ref}$ is integrated as a new state of the dynamical system using the same input constraints.
 The corresponding vector field $\dot{x}=f(x)+G(x)u_{sos}(x)$ (projected to $(\tilde{v}_{dc},\tilde{i}_{d})$ coordinates) is shown in blue, where $u_{sos}(x)=p(x)/s(x)$ is used to ensure compatibility between the CBF and CLF.
@@
 In this subsection, we assume a zero quadrature current reference, i.e., $i_{q,ref}=0$.
-The allowable set $\mathcal{X}_{a}$, as expressed in 12, encoding the state constraints 50, are achieved through the following polynomials:
-$w_{1}(x)$ $=((\tilde{v}_{dc}+0.3)/0.5)^{2}+(\tilde{i}_{d}/20)^{2}+(\tilde{i}_{q}/20)^{2}-1$
-$w_{2}(x)$ $=(\tilde{v}_{dc}/20)^{2}+({\tilde{i}_{d}}/1.3)^{2}+({\tilde{i}_{q}}/1.3)^{2}-1.$
-Figures 6 (a) and (b) illustrates the safe set $\mathcal{X}_{s}$ and the nominal set $\mathcal{X}_{n}$ computed by the algorithm both without and with considering inputs constraints.
+The allowable set $\mathcal{X}_{a}$, as expressed in 13, encoding the state constraints 60, are achieved through the following polynomials:
+$w_{1}(x)$ $=((\tilde{v}_{dc}+0.3)/0.5)^{2}-1$
+$w_{2}(x)$ $=({\tilde{i}_{d}}/1.3)^{2}+({\tilde{i}_{q}}/1.3)^{2}-1.$
+Figures 6 (a) and (b) illustrates the safe set $\mathcal{X}_{s}$ and the nominal region $\mathcal{X}_{n}$ computed by the algorithm both without and with considering inputs constraints.
 Table I summarizes both the number of variables associated with the respective SDPs and the number of constraints, derived from the summation of triangular elements within the positive semidefinite matrices.
@@
 The degree of the polynomials are given in Table II.
+Simulation results comparing the advanced safety filter with a basic safety filter are provided in the supplementary material.
 TABLE I: Summary of SDP variables and constraints
@@
 4 4 3 2
-The proposed safety filter introduces additional guarantees, which make it more stringent and, as noted in the introduction, more suitable in already established industrial setups.
-However, due these system requirements, it now represents a new category that prioritizes the preservation of the legacy control, enhancing composability, preventing a direct performance comparison between conventional safety filters and our novel setup.
-Nonetheless, side-by-side plots in Figure 7 are included that illustrate the additional guarantees compared to a basic safety filter in selected scenarios.
-The advanced safety filter $u_{s}(x)$ is based on the previously computed CLF $V(x)$ and CBFs $B_{1}(x)$ and $B_{2}(x)$, and is compared to a basic safety filter $u_{\text{SF}}(x)$ using the same CBFs:
-$\begin{array}[]{lll}\underset{u\in\mathcal{U}}{\mbox{argmin}}&\|u_{n}(x)-u\|^{2}\\ \mbox{s.t.}&\nabla B_{1}(x)(f(x)+G(x)u)+\alpha\left(B_{1}(x)\right)\leq 0\\ &\nabla B_{2}(x)(f(x)+G(x)u)+\alpha\left(B_{2}(x)\right)\leq 0,\end{array}$ (57)
-where $\alpha(B_{i}(x)):=10B_{i}(x)$ is selected.
-Figure 7: (Convergence to the nominal region $\mathcal{X}_{n}$): While the advanced safety filter $u_{s}(x)$ essentially overlaps the legacy control $u_{n}$, the basic safety filter $u_{\text{SF}}(x)$ suboptimally perturbs the legacy control action, resulting in a much slower convergence.
-Moreover, a side-by-side plots in Figure 8 illustrate the effect on including quadratic input constraints into the design of the safety filter.
-We consider the integration of quadratic input constraints into the SOS constraint as a contribution distinct from the discussion on the safety filter’s role in preserving the legacy control action.
-However, for consistency, the simulations were performed using the previous setup, which including an advanced safety filter and the basic safety filter in 57.
-Since the input constraints are incorporated into the design of the CLF and CBFs, the safety filter $u_{s}(x)$ successfully stays within the state constraints for all $t\geq 0$.
-This is not true for a basic safety filter $u_{\text{SF}}(x)$, which does not consider the input constraints in its design.
-In this case, the norm of the input action $\|u_{\text{SF}}(x)\|$ exceeds the input constraint limit of $1.3$ p.u.
-in the beginning of the simulation (cf.
-Figure 8).
-To produce a valid input, the input action $u_{\text{SF}}(x)$ is projected to the set of feasible input set $\mathcal{U}$.
-As a result, the trajectory deviates from its intended path, leading to a violation of the dc-link voltage constraint.
-Figure 8: (Forward invariance of the safe set $\mathcal{X}_{s}$): While the advanced safety filter $u_{s}(x)$, which considers the quadratic input constraints into the design, successfully stays within the state constraints, whereas the legacy controller $u_{n}$ and basic safety filter $u_{\text{SF}}(x)$ violate the state constraints.
 VI-B Including current reference tracking
-For this subsection, the objective is to stabilize both the DC-link voltage to $1$, and the quadrature current to some (not yet a priori pre-specified) quadrature current reference $i_{q,ref}$.
+For this subsection, the objective is to stabilize both the DC-link voltage to 11, and the quadrature current to some (not yet a priori pre-specified) quadrature current reference $i_{q,ref}$.
 To render our design parametric in $i_{q,ref}$, we introduce a new state representing the quadrature current reference.
-Consequently, we augment the dynamics from 49 with $\dot{i}_{q,ref}=0$.
+Consequently, we augment the dynamics from 59 with $\dot{i}_{q,ref}=0$.
 Remark 6
 Although the reference $i_{q,ref}$ is assumed constant for the design of the CBFs and CLF, step-wise adjustments to $i_{q,ref}$ over time (corresponding to task cycles of the digital processing unit) do not compromise safety, provided the state remains within the safe set $\mathcal{X}_{s}$.
-The corresponding coordinates of the states and inputs are given by $x=\begin{bmatrix}\tilde{v}_{dc}&\tilde{i}_{d}&\tilde{i}_{q}&\tilde{i}_{q,ref}\end{bmatrix}^{T}$ and $u:=\begin{bmatrix}\tilde{m}_{d}&\tilde{m}_{q}\end{bmatrix}^{T}$, where $\tilde{i}_{q,ref}=i_{q,ref}$.
-The allowable set $\mathcal{X}_{a}$ is chosen through the following polynomials:
-$w_{1}(x)$ $=((\tilde{v}_{dc}+0.3)/0.5)^{2}+(\tilde{i}_{d}/20)^{2}+(\tilde{i}_{q}/20)^{2}$
-$+(\tilde{i}_{q,ref}/20)^{2}-1$
-$w_{2}(x)$ $=(\tilde{v}_{dc}/20)^{2}+({\tilde{i}_{d}}/1.3)^{2}+({\tilde{i}_{q}}/1.3)^{2}$
-$+(\tilde{i}_{q,ref}/20)^{2}-1.$
+The corresponding coordinates of the states and inputs are given by $x=\begin{bmatrix}\tilde{v}_{dc}&\tilde{i}_{d}&\tilde{i}_{q}&\tilde{i}_{q,ref}\end{bmatrix}^{\top}$ and $u:=\begin{bmatrix}\tilde{m}_{d}&\tilde{m}_{q}\end{bmatrix}^{\top}$, where $\tilde{i}_{q,ref}=i_{q,ref}$.
 Correspondingly, a safe set $\mathcal{X}_{s}$ and the nominal region $\mathcal{X}_{n}$ can be computed for this 4-dimensional space.
-Figure 6(c) illustrates the safe set $\mathcal{X}_{s}$ and the nominal set $\mathcal{X}_{n}$ for $i_{q}=i_{q,ref}=0$.
+Figure 6(c) illustrates the safe set $\mathcal{X}_{s}$ and the nominal region $\mathcal{X}_{n}$ for $i_{q}=i_{q,ref}=0$.
 Table I summarizes the number of variables of the corresponding SDP for each part, as well as the number of constraints.
@@
 The main algorithm completed in 20 iterations, where each iteration took around 90 seconds.
-The degree of the polynomials are summarized in Table II.
-The simulation shown in Figure 9 illustrates the reference tracking performance of our safety filter $u_{s}(x)$ compared to the basic safety filter $u_{\text{SF}}(x)$ introduced in the previous subsection.
-The safety filter, although slightly perturbing the legacy controller, guarantees finite-time convergence to the nominal set defined for a quadrature current reference $i_{q,ref}$ of $1$.
-However, the basic safety filter $u_{\text{SF}}(x)$ perturbs the legacy control action, leading to a much slower convergence.
-In general, as we approach the boundary of the safe set by increasing $i_{q,ref}$, the basic safety filter is more likely to intervene the legacy control action.
-Figure 9: (Quadrature current reference tracking): The closed-loop simulation demonstrates the finite-time convergence guarantee of the advanced safety filter $u_{s}(x)$ towards the nominal region $\mathcal{X}_{n}$ for non-zero quadrature current reference $i_{q,ref}=1$p.u.
-Additionally, the advanced safety filter exhibits a faster convergence to the equilibrium compared to the basic safety filter $u_{\text{SF}}(x)$.
-The latter is more likely to disturb the legacy control action as the quadrature current reference $i_{q,ref}$ approaches the boundary of the safe set, resulting in a slower convergence.
 VII Conclusion
 We successfully developed and implemented a novel safety filter approach that, to our knowledge, is the first of its kind to leave the existing legacy controller subject to input constraints unaffected.
-This innovation could lead to greater acceptance of safety filters in industrial environments and real-world scenarios, advancing our understanding of their implementation across a broader range of applications.
-Our next step is to integrate this safety filter with our industrial partners for the control of an advanced converter topology.
+This innovation could lead to greater acceptance of safety filters in industrial environments and real-world scenarios, advancing their implementation across a broader range of applications.
+Unlike traditional safety filters, which often require manual tuning and lack formal guarantees on preserving the nominal controller, our approach provides an automated and principled alternative that ensures interpretability and minimal intervention by design.
+To support practical deployment, we also developed a software framework that automates the computation of the proposed safety filter.
+Future work will focus on a rigorous analysis of robustness to approximation errors in the legacy controller.
 VIII Appendix
@@
 Because the safe set $\mathcal{X}_{s}$ is compact and forward invariant, and the closed-loop dynamical system $\dot{x}=f(x)+G(x)u_{s}(x)$ is Lipschitz-continuous, there exists a unique trajectory $x(t)$ for all $x(0)\in\mathcal{X}_{s}$ and $t>0$.
-The CLF condition in 17 with $u:=u_{s}(x)$ can be written without dissipation rate as $\forall x\in\partial X_{n}$:
-$\nabla V(x)^{T}\left(f(x)+G(x)u_{s}(x)\right)\leq 0.$ (58)
-Using the tangent cone $T_{\mathcal{X}_{n}}(x)=\{z\mid\nabla V(x)^{T}z\leq 0\}$, the condition 58 can be formulated as $\forall x\in\partial X_{n}$:
+The CLF condition in 18 with $u:=u_{s}(x)$ can be written without dissipation rate as $\forall x\in\partial X_{n}$:
+$\nabla V(x)^{\top}\left(f(x)+G(x)u_{s}(x)\right)\leq 0.$ (63)
+Using the tangent cone $T_{\mathcal{X}_{n}}(x)=\{z\mid\nabla V(x)^{\top}z\leq 0\}$, the condition 63 can be formulated as $\forall x\in\partial X_{n}$:
 $f(x)+G(x)u_{s}(x)\in\mathcal{T}_{\mathcal{X}_{n}}(x).$
-Hence, according to [18, Nagumo’s Theorem 4.7], $\mathcal{X}_{n}$ is forward invariant.
+Hence, according to [5, Nagumo’s Theorem 4.7], $\mathcal{X}_{n}$ is forward invariant.
 This proves that for all $x(0)\in\mathcal{X}_{n}$, we can choose $T=0$.
@@
 As $\mathcal{X}_{t}\subseteq\mathcal{X}_{s}$ is compact (cf.
-Assumption 1) and the dissipative rate $d(x)$ defined in 18 is assumed to be strictly positive and Lipschitz-continuous on $\mathcal{X}_{t}$, the following minimum exists: $l=\underset{x\in\mathcal{X}_{t}}{\text{min}}\lambda(x)\left(V(x)+V_{0}\right)>0$.
-Therefore, the scalar function $V\left(x(t)\right)$ strictly decreases over time for all $x(0)\in\mathcal{X}_{t}$ due to the CLF condition in 17: $\dot{V}=\nabla V(x)\left(f(x)+G(x)u_{s}(x)\right)\leq-l<0$.
+Assumption 1) and the dissipative rate $d(x)$ defined in 19 is assumed to be strictly positive and Lipschitz-continuous on $\mathcal{X}_{t}$, the following minimum exists: $l=\underset{x\in\mathcal{X}_{t}}{\text{min}}\lambda(x)\left(V(x)+V_{0}\right)>0$.
+Therefore, the scalar function $V\left(x(t)\right)$ strictly decreases over time for all $x(0)\in\mathcal{X}_{t}$ due to the CLF condition in 18: $\dot{V}=\nabla V(x)\left(f(x)+G(x)u_{s}(x)\right)\leq-l<0$.
 As $\mathcal{X}_{s}$ is forward invariant, $V\left(x(t)\right)$ is upper bounded by $V\left(x(t)\right)\leq V\left(x(0)\right)-lt$ for all $x(0)\in\mathcal{X}_{t}$.
 Therefore, there exists some $T\leq V\left(x(0)\right)/l$ such that $V\left(x(T)\right)\leq 0$.
-VIII-B Proof of Lemma 2
-By dividing 26a and 26b by $s(x)>0$, we obtain
-$\begin{split}\nabla B_{i}^{T}\left(f+Gu_{\text{sos}}\right)+\frac{\gamma_{0,i}}{s}B_{i}-\frac{\gamma_{1,i}^{T}}{s}B_{\mathcal{I}_{t}}&\leq 0\end{split}$ (59a)
-$\begin{split}\nabla V^{T}\left(f+Gu_{\text{sos}}\right)+d+\frac{\gamma_{0,0}}{s}V-\frac{\gamma_{1,0}^{T}}{s}B_{\mathcal{I}_{t}}&\leq 0,\end{split}$ (59b)
+VIII-B Derivation of the SOS constraints 26
+The objective of this section is to convert the empty-set condition 24b into the SOS constraint given in 26b.
+The derivation of 26a and 26c is analogous.
+Applying Putinar’s P-satz to 24b yields the SOS constraint
+$-\nabla V^{T}f-d\in\textbf{M}(\nabla V^{T}G,-\nabla V^{T}G,V,-B_{\mathcal{I}_{t}}).$
+By utilizing the definition of a quadratic module in 5, we can express this SOS constraint by the condition that there exist two SOS polynomials $p_{0,1},p_{0,2}\in(\Sigma[x])^{m}$ such that
+$-\nabla V^{T}\left(f+G(p_{0,1}-p_{0,2})\right)-d\in\textbf{M}(V,-B_{\mathcal{I}_{t}}).$
+By exploiting the fact that the set of SOS polynomials, when combined with the set of negative SOS polynomials, spans the entire set of polynomials (i.e., $\Sigma[x]-\Sigma[x]=R[x]$), we can reformulate the SOS constraint as the condition that there exists a polynomial $p_{0}\in(R[x])^{m}$ such that
+$-\nabla V^{T}\left(f+Gp_{0}\right)-d\in\textbf{M}(V,-B_{\mathcal{I}_{t}}).$
+Finally, we introduce an additional SOS polynomial $s_{0}(x)$ (cf. 8) and rewrite the SOS constraint as the condition that there exists $s_{0}\in\Sigma[x],p_{0}\in(R[x])^{m}$ such that
+$-\nabla V^{T}\left(s_{0}f+Gp_{0}\right)-s_{0}d\in\textbf{M}(V,-B_{\mathcal{I}_{t}}).$
+The introduction of the polynomial $s_{0}(x)$ increases the set of feasible solutions when truncating the degrees of the involved polynomials, as described in Section II-B.
+VIII-C Proof of Lemma 2
+By dividing 27a and 27b by $s(x)>0$, we obtain
+$\begin{split}\nabla B_{i}^{\top}\left(f+Gu_{\text{sos}}\right)+\frac{\gamma_{0,i}}{s}B_{i}-\frac{\gamma_{1,i}^{\top}}{s}B_{\mathcal{I}_{n_{\!B}}}&\leq 0\end{split}$ (64a)
+$\begin{split}\nabla V^{\top}\left(f+Gu_{\text{sos}}\right)+d+\frac{\gamma_{0,0}}{s}V-\frac{\gamma_{1,0}^{\top}}{s}B_{\mathcal{I}_{n_{\!B}}}&\leq 0,\end{split}$ (64b)
 where $u_{\text{sos}}(x)=p(x)/s(x)$.
-From 59a, by using the fact that $\gamma_{1,i}(x)\geq 0$, $s(x)>0$, and $B_{i}(x)=0$ and $B_{\mathcal{I}_{t}}(x)\leq 0$ on $\partial\mathcal{X}_{s}$, we infer, for each $i\in\mathcal{I}_{t}$, that for all $x\in\partial\mathcal{X}_{s}$ there exists $u:=u_{\text{sos}}(x)$ such that $\nabla B_{i}^{T}\left(f+Gu_{\text{sos}}\right)\leq 0$ whenever $B_{i}(x)=0$.
-Or, equivalently, for all $x\in\partial\mathcal{X}_{s}$ there exists $u:=u_{\text{sos}}(x)$ such that $\nabla B_{i}^{T}\left(f+Gu\right)\leq 0$ for all $i\in\text{Act}(x)$, as defined in 16.
-This proofs the CBF condition in 15 using the definition of the tangent cone under Assumption 2.
-From 59b, by using the fact that $\gamma_{0,0}(x)\geq 0$, $\gamma_{1,0}(x)\geq 0$, $s(x)>0$, and $V(x)\geq 0$ and $B_{\mathcal{I}_{t}}(x)\leq 0$ on $\mathcal{X}_{t}$, we deduce that for all $x\in\mathcal{X}_{t}$ there exists $u:=u_{\text{sos}}(x)$ such that $\nabla V^{T}\left(f+Gu\right)+\lambda\bigl{(}V-V_{0}\bigr{)}\leq 0$.
-This proofs the CLF condition 17.
-Finally, from 26c, by using the fact that $V(x)=0$ on $\partial\mathcal{X}_{n}$, we deduce that the inequality $\nabla V^{T}\left(f+Gu_{n}\right)\leq 0$ holds for all $x\in\partial\mathcal{X}_{n}$.
-This proofs the CBF condition in 19.
-VIII-C Proof of Lemma 3
-We use the definition of a quadratic module in 4, to rewrite the SOS constraint in 28 as $\forall x\in\mathbb{R}^{n},\exists\gamma_{u}\in(\Sigma[x])^{t}$:
-$-a_{i}u_{\text{SOS}}(x)+b_{i}+\frac{\gamma_{u}(x)^{T}}{s(x)}B_{\mathcal{I}_{t}}(x)\geq 0$
+From 64a, by using the fact that $\gamma_{1,i}(x)\geq 0$, $s(x)>0$, and $B_{i}(x)=0$ and $B_{\mathcal{I}_{n_{\!B}}}(x)\leq 0$ on $\partial\mathcal{X}_{s}$, we infer, for each $i\in\mathcal{I}_{n_{\!B}}$, that for all $x\in\partial\mathcal{X}_{s}$ there exists $u:=u_{\text{sos}}(x)$ such that $\nabla B_{i}^{\top}\left(f+Gu_{\text{sos}}\right)\leq 0$ whenever $B_{i}(x)=0$.
+Or, equivalently, for all $x\in\partial\mathcal{X}_{s}$ there exists $u:=u_{\text{sos}}(x)$ such that $\nabla B_{i}^{\top}\left(f+Gu\right)\leq 0$ for all $i\in\text{Act}(x)$, as defined in 17.
+This proofs the CBF condition in 16 using the definition of the tangent cone under Assumption 2.
+From 64b, by using the fact that $\gamma_{0,0}(x)\geq 0$, $\gamma_{1,0}(x)\geq 0$, $s(x)>0$, and $V(x)\geq 0$ and $B_{\mathcal{I}_{n_{\!B}}}(x)\leq 0$ on $\mathcal{X}_{t}$, we deduce that for all $x\in\mathcal{X}_{t}$ there exists $u:=u_{\text{sos}}(x)$ such that $\nabla V^{\top}\left(f+Gu\right)+\lambda\bigl(V-V_{0}\bigr)\leq 0$.
+This proofs the CLF condition 18.
+Finally, from 27c, by using the fact that $V(x)=0$ on $\partial\mathcal{X}_{n}$, we deduce that the inequality $\nabla V^{\top}\left(f+Gu_{n}\right)\leq 0$ holds for all $x\in\partial\mathcal{X}_{n}$.
+This proofs the CBF condition in 20.
+VIII-D Proof of Lemma 3
+We use the definition of a quadratic module in 5, to rewrite the SOS constraint in 29 as $\forall x\in\mathbb{R}^{n},\exists\gamma_{u}\in(\Sigma[x])^{\top}$:
+$-a_{i}u_{\text{SOS}}(x)+b_{i}+\frac{\gamma_{u}(x)^{\top}}{s(x)}B_{\mathcal{I}_{n_{\!B}}}(x)\geq 0$
 for $i\in\mathcal{I}_{l}$, and $u_{\text{SOS}}(x)=p(x)/s(x)$.
-Since the $\gamma_{u}(x)\geq 0$, $s(x)>0$, and $B_{\mathcal{I}_{t}}(x)\leq 0$ on $\mathcal{X}_{s}$, we deduce, for all $i\in\mathcal{I}_{l}$, $-a_{i}u_{\text{SOS}}(x)+b_{i}\geq 0$ on $\mathcal{X}_{s}$.
+Since the $\gamma_{u}(x)\geq 0$, $s(x)>0$, and $B_{\mathcal{I}_{n_{\!B}}}(x)\leq 0$ on $\mathcal{X}_{s}$, we deduce, for all $i\in\mathcal{I}_{l}$, $a_{i}u_{\text{SOS}}(x)+b_{i}\geq 0$ on $\mathcal{X}_{s}$.
 This concludes the first part of the proof.
-By substituting the variable $v$ in 30 with $u-u_{0}$, we can reformulate the SOS constraints by using the definition of a quadratic module as $\forall x\in\mathbb{R}^{n},\exists\gamma_{u,1}\in\Sigma[x,u],\gamma_{u,2}\in(\Sigma[x,u])^{t}$:
-$\begin{split}&-\left((u-u_{0})^{T}Q_{u}\left(u_{\text{SOS}}(x)-u_{0}\right)-u_{\text{max}}^{2}\right)\\ &\quad+\frac{\gamma_{u,1}(x,u)}{s(x)}f_{u}(x,u-u_{0})+\frac{\gamma_{u,2}(x,u)^{T}}{s(x)}B_{\mathcal{I}_{t}}\geq 0,\end{split}$ (60)
+By substituting the variable $v$ in 31 with $u-u_{0}$, we can reformulate the SOS constraints by using the definition of a quadratic module as $\forall x\in\mathbb{R}^{n},\exists\gamma_{u,1}\in\Sigma[x,u],\gamma_{u,2}\in(\Sigma[x,u])^{\top}$:
+$\begin{split}&-\left((u-u_{0})^{\top}Q_{u}\left(u_{\text{SOS}}(x)-u_{0}\right)-u_{\text{max}}^{2}\right)\\ &\quad+\frac{\gamma_{u,1}(x,u)}{s(x)}f_{u}(x,u-u_{0})+\frac{\gamma_{u,2}(x,u)^{\top}}{s(x)}B_{\mathcal{I}_{n_{\!B}}}\geq 0,\end{split}$ (65)
 where $x\in\mathbb{R}^{n}$ and $u\in\mathbb{R}^{m}$ are the variables of the polynomial expressions.
-Because of its positive semidefiniteness, $Q_{u}$ can be decomposed into $Q_{u}=B_{u}^{T}B_{u}$.
-Furthermore, since $\gamma_{u,1}(x,u)\geq 0$, $\gamma_{u,2}(x,u)\geq 0$, $s(x)>0$, $B_{\mathcal{I}_{t}}(x)\leq 0$ on $\mathcal{X}_{s}$, and $f_{u}(x,u-u_{0})\leq 0$ on $\mathcal{U}$, condition 60 can be reformulated as $\forall x\in\mathcal{X}_{s},\forall u\in\mathcal{U}$:
-$(u-u_{0})^{T}B_{u}^{T}B_{u}(u_{\text{SOS}}(x)-u_{0})\leq u_{\text{max}}^{2}.$ (61)
-For any $x\in\{x\in\mathcal{X}_{s}\mid B_{u}(u_{\text{SOS}}(x)-u_{0})=0\}$, we have $(u_{\text{SOS}}(x)-u_{0})^{T}B_{u}^{T}B_{u}(u_{\text{SOS}}(x)-u_{0})=0\leq u_{\text{max}}^{2}$.
+Because of its positive semidefiniteness, $Q_{u}$ can be decomposed into $Q_{u}=B_{u}^{\top}B_{u}$.
+Furthermore, since $\gamma_{u,1}(x,u)\geq 0$, $\gamma_{u,2}(x,u)\geq 0$, $s(x)>0$, $B_{\mathcal{I}_{n_{\!B}}}(x)\leq 0$ on $\mathcal{X}_{s}$, and $f_{u}(x,u-u_{0})\leq 0$ on $\mathcal{U}$, condition 65 can be reformulated as $\forall x\in\mathcal{X}_{s},\forall u\in\mathcal{U}$:
+$(u-u_{0})^{\top}B_{u}^{\top}B_{u}(u_{\text{SOS}}(x)-u_{0})\leq u_{\text{max}}^{2}.$ (66)
+For any $x\in\{x\in\mathcal{X}_{s}\mid B_{u}(u_{\text{SOS}}(x)-u_{0})=0\}$, we have $(u_{\text{SOS}}(x)-u_{0})^{\top}B_{u}^{\top}B_{u}(u_{\text{SOS}}(x)-u_{0})=0\leq u_{\text{max}}^{2}$.
 Hence, $u_{\text{SOS}}(x)\in\mathcal{U}$.
 For any other safe state, that is $x\in\{x\in\mathcal{X}_{s}\mid B_{u}(u_{\text{SOS}}(x)-u_{0})\neq 0\}$, there exists a unique $\lambda>0$ such that
-$\begin{split}\left(B_{u}\lambda(u_{\text{SOS}}(x)-u_{0})\right)^{T}\left(B_{u}\lambda(u_{\text{SOS}}(x)-u_{0})\right)&=u_{\text{max}}^{2}.\end{split}$ (62)
-Additionally, substituting $u$ by $\lambda(u_{\text{SOS}}(x)-u_{0})+u_{0}$ in condition 61 yields the condition that $\forall x\in\mathcal{X}_{s}$:
-$\lambda(B_{u}(u_{\text{SOS}}(x)-u_{0}))^{T}(B_{u}(u_{\text{SOS}}(x)-u_{0}))\leq u_{\text{max}}^{2}.$ (63)
-By comparing 62 and 63, we conclude that $\lambda$ must be greater than $1$.
-As a consequence, $(B_{u}(u_{\text{SOS}}(x)-u_{0}))^{T}(B_{u}(u_{\text{SOS}}(x)-u_{0}))\leq u_{\text{max}}^{2}$ holds for all $x\in\mathcal{X}_{s}$.
+$\begin{split}\left(B_{u}\lambda(u_{\text{SOS}}(x)-u_{0})\right)^{\top}\left(B_{u}\lambda(u_{\text{SOS}}(x)-u_{0})\right)&=u_{\text{max}}^{2}.\end{split}$ (67)
+Additionally, substituting $u$ by $\lambda(u_{\text{SOS}}(x)-u_{0})+u_{0}$ in condition 66 yields the condition that $\forall x\in\mathcal{X}_{s}$:
+$\lambda(B_{u}(u_{\text{SOS}}(x)-u_{0}))^{\top}(B_{u}(u_{\text{SOS}}(x)-u_{0}))\leq u_{\text{max}}^{2}.$ (68)
+By comparing 67 and 68, we conclude that $\lambda$ must be greater than 11.
+As a consequence, $(B_{u}(u_{\text{SOS}}(x)-u_{0}))^{\top}(B_{u}(u_{\text{SOS}}(x)-u_{0}))\leq u_{\text{max}}^{2}$ holds for all $x\in\mathcal{X}_{s}$.
 This concludes that $u_{\text{SOS}}(x)\in\mathcal{U}$ for all $x\in\mathcal{X}_{s}$.
-VIII-D Proof of Theorem 2
+VIII-E Proof of Theorem 2
 Consider the optimization problem
@@
 We say $(u_{s}(x^{*}),\lambda(x^{*}))$ is a regular minimizer if the LICQ is satisfied at $u_{s}(x^{*})$, and $(u_{s}(x^{*}),\lambda(x^{*}))$ satisfies the KKT condition and the strong second order sufficient condition (SSOSC).
-The point $(u,\lambda)$ satisfies the SSOSC if Hessian matrix of the Lagrangian of 66 is positive definite at this point.
-According to [25, Theorem 2.3.3], there exists a locally Lipschitz continuous map $u_{s}(x)$ around a neighborhood of $x^{*}$ such that $u_{s}(x)\in\mathbb{R}^{m}$ is a minimizer for the non-linear optimization problem 66, if $f$ and $g$ are twice continuously differentiable in $(u,x)$, and $u_{s}(x^{*})$ is a regular minimizer with multiplier $\lambda(x^{*})$.
-In what follows, we establish these conditions for the QCQP in 43.
-LICQ property under Assumption 3 asserts that for any solution of the QCQP $u_{s}(x^{*})$, there exists a unique $\lambda^{*}$ that satisfies the KKT condition at $(u_{s}(x^{*}),\lambda^{*})$.
+The point $(u,\lambda)$ satisfies the SSOSC if Hessian matrix of the Lagrangian of 71 is positive definite at this point.
+According to [9, Theorem 2.3.3], there exists a locally Lipschitz continuous map $u_{s}(x)$ around a neighborhood of $x^{*}$ such that $u_{s}(x)\in\mathbb{R}^{m}$ is a minimizer for the non-linear optimization problem 71, if $f$ and $g$ are twice continuously differentiable in $(u,x)$, and $u_{s}(x^{*})$ is a regular minimizer with multiplier $\lambda(x^{*})$.
+In what follows, we establish these conditions for the QP in 44.
+LICQ property under Assumption 3 asserts that for any solution of the QP $u_{s}(x^{*})$, there exists a unique $\lambda^{*}$ that satisfies the KKT condition at $(u_{s}(x^{*}),\lambda^{*})$.
 Next, we show that the SSOSC holds for any choice of $\lambda^{*}$, specifically $\lambda^{*}$ that satisfies the KKT condition.
-Consider the Lagrangian of the QCQP for either linear or quadratic input constraints:
-$\begin{split}L(u,\lambda)=&u^{T}u+2u_{n}(x)u+u_{n}(x)^{T}u_{n}(x)\\ &+\lambda_{1}^{T}(C(x)u+b(x))\\ &+\lambda_{2}^{T}(u^{T}Q_{u}u-u_{max}^{2})+\lambda_{3}^{T}(Au-b)\end{split}$
-Then the inequality $v^{T}\nabla_{uu}^{2}L(u,\lambda)v=v^{T}v+\lambda_{2}v^{T}Q_{u}v>0$ for all $v\in\mathbb{R}^{m}$.
-This proves that any solution of the QCQP $u_{s}(x^{*})$ satisfies the SSOSC with the corresponding $\lambda^{*}$.
-Furthermore, a solution of the QCQP $u_{s}(x)$ exists for all $x$ due to the lower bound on $r(x)$ in 48, consequently establishing the local Lipschitz continuity of $u_{s}(x)$.
+Consider the Lagrangian of the QP for either linear or quadratic input constraints:
+$\begin{split}L(u,\lambda)=&u^{\top}u+2u_{n}(x)u+u_{n}(x)^{\top}u_{n}(x)\\ &+\lambda_{1}^{\top}(C(x)u+b(x))\\ &+\lambda_{2}^{\top}(u^{\top}Q_{u}u-u_{max}^{2})+\lambda_{3}^{\top}(Au-b)\end{split}$
+Then the inequality $v^{\top}\nabla_{uu}^{2}L(u,\lambda)v=v^{\top}v+\lambda_{2}v^{\top}Q_{u}v>0$ for all $v\in\mathbb{R}^{m}$.
+This proves that any solution of the QP $u_{s}(x^{*})$ satisfies the SSOSC with the corresponding $\lambda^{*}$.
+Furthermore, a solution of the QP $u_{s}(x)$ exists for all $x$ due to the lower bound on $r(x)$ in 49, consequently establishing the local Lipschitz continuity of $u_{s}(x)$.